### 12 Jan 2024 v1

# A spectral approach to the narrow escape problem in the disk

Abstract.
We study the narrow escape problem in the disk, which consists in identifying the first exit time and first exit point distribution of a Brownian particle from the ball in dimension 2, with reflecting boundary conditions except on small disjoint windows through which it can escape. This problem is motivated by practical questions arising in various scientific fields (cellular biology, molecular dynamics). We apply the quasi-stationary distribution approach to metastability, which requires to study the eigenvalue problem for the Laplacian operator with Dirichlet boundary conditions on the small absorbing part of the boundary, and Neumann boundary conditions on the remaining reflecting part. We obtain rigorous asymptotic expansions of the first eigenvalue and of the normal derivative of the associated eigenfunction in the limit of infinitely small exit regions, which yields asymptotic expansions of the first exit time and first exit point distribution starting from the quasi-stationary distribution within the disk.
1. Introduction and Motivation
The narrow escape problem is a question arising in various models used in biophysics and cell biology. The mathematical formulation of the problem is as follows: a Brownian particle (representing e.g. an ion or a protein) is confined to a bounded domain (representing e.g. a biological cell), with a reflecting boundary except for small disjoint windows (representing e.g. membrane channels) through which it can escape. The narrow escape problem then consists in precisely describing the exit event, namely the law of the first exit time and the distribution of the first exit point. One is for example interested in the average time needed for an ion to find an open ion channel located in the cell membrane, and also in information on which exit channel will be the most likely. Because of the practical importance of this question, there are numerous contributions from the physics community, which concentrate on the asymptotic behavior of the exit time, and mostly on the mean first exit time (see however [18] for studies on the law of the exit time). We refer for example to [24, 5, 19, 46] for a few representative contributions from the physics literature, and for typical practical applications in biology. Our objective in this work is to prove asymptotic results on both the first exit time distribution and the first exit point distribution, using the quasi-stationary distribution approach to metastability [11].
1.1. Mathematical framework
Let us introduce the mathematical model associated with the problem. The motion of a particle in a bounded domain $\Omega\subset\mathbb{R}^{d}$ is described by the diffusion process:
(1.1) $dX_{t}=\sqrt{2}\hskip 1.42271ptdB_{t},$
where $X_{t}\in\Omega$ and $(B_{t})_{t\geqslant 0}$ is a $d$-dimensional Brownian motion. Here $\Omega\subset\mathbb{R}^{d}$ is a bounded domain whose boundary is $\partial\Omega=\overline{\Gamma^{\varepsilon}_{\mathbf{D}}}\cup\overline{\Gamma^{\varepsilon}_{\mathbf{N}}}$ with $\Gamma^{\varepsilon}_{\mathbf{D}}\cap\Gamma^{\varepsilon}_{\mathbf{N}}=\emptyset$. The set $\Gamma^{\varepsilon}_{\mathbf{D}}={\cup}_{k=1}^{N}\Gamma^{\varepsilon}_{\mathbf{D}_{k}}$ is composed of $N$ small disjoint connected absorbing windows $(\Gamma^{\varepsilon}_{\mathbf{D}_{k}})_{k=1,\ldots,N}$, and $\Gamma^{\varepsilon}_{\mathbf{N}}$ is the reflecting part (see Figure 1 for a schematic representation in dimension $d=2$). More precisely, let us introduce $N$ points $x^{(1)},\dots,x^{(N)}$ on the boundary of $\Omega$ and let $\rho_{0}\in(0,2]$ be defined as
(1.2) $\rho_{0}=\min_{1\leqslant i<j\leqslant N}\left|x^{(i)}-x^{(j)}\right|.$
We then assume that [^1] The main results (Theorems 1.3 and 1.4) will actually be obtained for $\Omega$ the unit disk (thus in dimension $d=2$), and a slightly different geometry for the escape regions that coincides with the description given here in the limit $\varepsilon\to 0$. This is made precise below. there exists $\varepsilon_{0}>0$ (which we will need to further reduce later on) such that for any $\varepsilon\in(0,\varepsilon_{0})$ and $k\in\{1,\dots,N\}$, there exists $K^{(k)}_{\varepsilon}>0$ satisfying
(1.3) $\Gamma^{\varepsilon}_{\mathbf{D}_{k}}=\partial\Omega\cap\mathbf{B}\left(x^{(k)},\mathrm{e}^{-1/K^{(k)}_{\varepsilon}}\right),\qquad\mathrm{e}^{-1/K^{(k)}_{\varepsilon}}\leqslant\frac{\rho_{0}}{2},$
where $\lim_{\varepsilon\to 0}K^{(k)}_{\varepsilon}=0$, and where $\mathbf{B}(x,r)$ denotes the disk with center $x$ and radius $r$. The second condition in (1.3) ensures that the holes are disjoint. To give a concrete example, $N$ identical holes with sizes of order $\varepsilon$ are obtained by considering $K^{(k)}_{\varepsilon}=1/|\log\varepsilon|$.
For $(X_{t})_{t\geqslant 0}$ satisfying (1.1) with an initial condition $X_{0}\in\Omega$, we define the first exit time of $(X_{t})_{t\geqslant 0}$ as follows:
$\tau:=\text{inf}\{t\geqslant 0,\hskip 2.84544ptX_{t}\notin\Omega\}.$
The exit event from $\Omega$ is fully characterized by the couple of random variables $(\tau,X_{\tau})$.
----||||||||----$x^{(2)}$$x^{(3)}$$x^{(1)}$$\Gamma^{\varepsilon}_{\mathbf{D}_{2}}$$\Gamma^{\varepsilon}_{\mathbf{D}_{1}}$$\Gamma^{\varepsilon}_{\mathbf{D}_{3}}$$\Gamma^{\varepsilon}_{\mathbf{N}}$$\Gamma^{\varepsilon}_{\mathbf{N}}$$\Gamma^{\varepsilon}_{\mathbf{N}}$$\Omega$$x_{0}$
Figure 1. Example of a domain $\Omega$, ${\Gamma}^{\varepsilon}_{\mathbf{N}}$ and ${\Gamma}^{\varepsilon}_{\mathbf{D}}$ (for $d=2$ and $N=4$).
Since the escape regions are very narrow, the process $(X_{t})_{t\geqslant 0}$ is metastable: it stays trapped for a very long time in the domain $\Omega$ before leaving it through one of the small windows. As a consequence, we use the quasi-stationary distribution approach to metastability: we study the exit event for the stochastic process $(X_{t})_{t\geqslant 0}$ starting from the quasi-stationary distribution within $\Omega$. We refer to [27, 11, 29, 33] for a justification of this approach. The quasi-stationary distribution can be defined as the distribution $\nu_{0}^{\varepsilon}$ such that, for any initial distribution $\mu$,
(1.4) $\lim_{t\to\infty}\mathbb{P}_{\mu}[X_{t}\in\cdot\,|\,t<\tau]=\nu_{0}^{\varepsilon},$
where the subscript $\mu$ in $\mathbb{P}_{\mu}$ indicates that $X_{0}\sim\mu$. In words, the quasi-distribution is the longtime limit of the distribution of configurations associated with trajectories which have not exited $\Omega$. The distribution $\nu_{0}^{\varepsilon}$ is thus uniquely defined, with support in $\overline{\Omega}$.
Our aim is to determine asymptotic estimates of the law of the first exit time $\tau$ and of the first exit point $X_{\tau}$ as $\varepsilon\rightarrow 0$, with $X_{0}\sim\nu_{0}^{\varepsilon}$. A crucial result which explains the interest of considering the exit event starting from $\nu_{0}^{\varepsilon}$ is the following: if $X_{0}\sim\nu_{0}^{\varepsilon}$, then $\tau$ is exponentially distributed and independent of $X_{\tau}$ (see [27, 9]). Let us mention that this result is in particular fundamental to justify the use of the kinetic Monte Carlo algorithm (see [52, 11]) to efficiently sample the exit event, and thus simulate the process over very long times. It therefore remains to identify $\mathbb{E}_{\nu_{0}^{\varepsilon}}[\tau]$ and the distribution of $X_{\tau}$ to fully characterize the law of $(\tau,X_{\tau})$. Here and in the following, the superscript in $\mathbb{E}_{\nu_{0}^{\varepsilon}}[\cdot]$ indicates that the process $(X_{t})_{t\geqslant 0}$ is such that $X_{0}\sim\nu_{0}^{\varepsilon}$.
Let us now explain why the above question actually amounts to studying the behavior of the first eigenvalue and first eigenvector of the infinitesimal generator of the stochastic process. The (opposite [^2] The sign convention is justified by the fact that we prefer to work with positive operators. of the) infinitesimal generator of the dynamics (1.1) is $-\Delta$. In particular, the density function $f(t,x)$ of the process $X_{t}$ reflected on $\Gamma^{\varepsilon}_{\mathbf{N}}$ and absorbed on $\Gamma^{\varepsilon}_{\mathbf{D}}$ satisfies the Fokker-Planck equation:
$\partial_{t}f&=\Delta f\text{ in }\Omega,\\ \partial_{n}f&=0\text{ on }\Gamma^{\varepsilon}_{\mathbf{N}}\text{ and }f=0\text{ on }\Gamma^{\varepsilon}_{\mathbf{D}},\\ f(0,\cdot)&=f_{0}\text{ in }\Omega,\end{aligned}\right.$
where $f_{0}$ is the density of $X_{0}$, and $\partial_{n}=\vec{n}\cdot\nabla$ denotes the normal derivative with $\vec{n}$ the unit vector outward normal to $\Omega$. Let us introduce the principal eigenvalue $\lambda_{0}^{\varepsilon}$ and the associated eigenfunction $u_{0}^{\varepsilon}$ related to this Fokker-Planck equation:
(1.5) $-\Delta u_{0}^{\varepsilon}&=\lambda_{0}^{\varepsilon}\,u_{0}^{\varepsilon}\,\,\text{in}\,\,\Omega,\\ \partial_{n}u_{0}^{\varepsilon}&=0\,\,\,\text{on}\,\,\Gamma^{\varepsilon}_{\mathbf{N}},\\ u_{0}^{\varepsilon}&=0\,\,\,\text{on}\,\,\Gamma^{\varepsilon}_{\mathbf{D}}.\end{aligned}\right.$
We denote by $\mathcal{L}^{\varepsilon}$ the operator [^3] The superscript $\varepsilon$ indicates that the boundary conditions, and thus the domain of the operator, depend on $\varepsilon$ (see the definition (1.12) below). associated with (1.5), namely minus the Laplacian with mixed boundary conditions on $\partial\Omega$. In the following, we normalize $u_{0}^{\varepsilon}$ as
(1.6) $\int_{\Omega}(u_{0}^{\varepsilon}(x))^{2}\,\mathrm{d}x=1,\qquad u_{0}^{\varepsilon}<0\textrm{ on }\Omega.$
The second condition can indeed be imposed as it is standard to show that the first eigenfunction $u^{\varepsilon}_{0}$ does not vanish in $\Omega$ (see for example [16, Theorem 8.38]), so that it can be chosen to be negative in $\Omega$.
Let us explain why studying the exit problem starting from the quasi-stationary distribution reduces to studying the principal eigenvalue and eigenfunction (see [27]). The unique quasi-stationary distribution $\nu_{0}^{\varepsilon}$ of the process (1.1) in $\Omega$ can be written as follows:
$\nu_{0}^{\varepsilon}(\mathrm{d}x)=\frac{u_{0}^{\varepsilon}(x)\,\mathrm{d}x}{\displaystyle\int_{\Omega}u_{0}^{\varepsilon}(y)\,\mathrm{d}y}.$
Moreover, starting from $\nu_{0}^{\varepsilon}$, the exit time $\tau$ is exponentially distributed with parameter $\lambda_{0}^{\varepsilon}$:
(1.7) $\mathbb{E}_{\nu_{0}^{\varepsilon}}[\tau]=\frac{1}{\lambda_{0}^{\varepsilon}}.$
In addition, the law of $X_{\tau}$ (with support on $\partial\Omega$) is given by
(1.8) $\frac{\partial_{n}u_{0}^{\varepsilon}(x)\,\sigma(dx)}{\displaystyle\int_{\partial\Omega}\partial_{n}u_{0}^{\varepsilon}(y)\,\sigma(\mathrm{d}y)},$
where $\sigma$ denotes generically in the following Lebesgue surface measures, here on $\partial\Omega$. In particular, the probability of exiting through the window $\Gamma^{\varepsilon}_{\mathbf{D}_{k}}$ is
(1.9) $\mathbb{P}_{\nu_{0}^{\varepsilon}}[X_{\tau}\in\Gamma^{\varepsilon}_{\mathbf{D}_{k}}]=\frac{\displaystyle\int_{\Gamma^{\varepsilon}_{\mathbf{D}_{k}}}\partial_{n}u_{0}^{\varepsilon}(x)\,\sigma(\mathrm{d}x)}{\displaystyle\int_{\Gamma^{\varepsilon}_{\mathbf{D}}}\partial_{n}u_{0}^{\varepsilon}(y)\,\sigma(\mathrm{d}y)}.$
In order to characterize the law of $(\tau,X_{\tau})$ in the limit $\varepsilon\rightarrow 0$, it thus suffices to study the asymptotic behavior of the first eigenvalue $\lambda_{0}^{\varepsilon}$ and of the normal derivative of $u_{0}^{\varepsilon}$ in this limiting regime.
In previous works [37, 51], such an approach has been used to study the exit event of the overdamped Langevin dynamics from the basin of attraction of a local minimum of a potential enegy function: in this context, metastability is related to energetic barriers that the process has to overcome in order to leave a metastable basin. In particular, the aforementioned works justify the use of the so-called Eyring–Kramers formulas to parameterize the law of the exit event. In the present work, we initiate a similar analysis for a situation where metastability is not due to energetic barriers but to entropic barriers: the process does not have to climb over energy levels to leave $\Omega$, but to find the narrow exit regions on the boundary. Let us emphasize that the present work should be seen as a first application of the quasi-stationary distribution approach to metastability originating from such purely entropic barriers. As discussed below, it opens the route to many generalizations (to other dynamics, and other geometries) and to numerical algorithms (to efficiently simulate the exit event relying on the identified asymptotic behaviours).
1.2. Main results.
Let us now present our main results. To perform our analysis, we consider the case when $\Omega$ is the unit disk, i.e.
$\Omega=\mathbf{B}(0,1).$
The first one concerns the spectral analysis of the operator $\mathcal{L}^{\varepsilon}$. A relevant parameter for the analysis is
(1.10) $\overline{K}_{\varepsilon}=\displaystyle\sum_{\begin{subarray}{c}k=1\end{subarray}}^{N}K^{(k)}_{\varepsilon}.$
Upon reducing $\varepsilon_{0}$, one can assume without loss of generality that, for all $\varepsilon\in(0,\varepsilon_{0})$,
(1.11) $\overline{K}_{\varepsilon}<\frac{1}{2}\min\left(\frac{1}{|\log(\rho_{0}/2)|},1\right).$
Theorem 1.1.
Let $\mathcal{L}^{\varepsilon}$ be the unbounded operator defined on $L^{2}(\Omega)$ with domain
(1.12) $\mathcal{D}(\mathcal{L}^{\varepsilon})=\left\{u\in H^{1}(\Omega),\,\,\Delta u\in L^{2}(\Omega),\,\,\partial_{n}u_{|_{\Gamma^{\varepsilon}_{\mathbf{N}}}}=0,\,\,u_{|_{\Gamma^{\varepsilon}_{\mathbf{D}}}}=0\right\}.$
Then,
  1. (i)
    the operator $\mathcal{L}^{\varepsilon}$ is nonnegative self-adjoint and has a compact resolvent;
  2. (ii)
    there exist $c>0$ and $\varepsilon_{0}>0$ such that, for any $\varepsilon\in(0,\varepsilon_{0})$,
    $\operatorname{dim}\operatorname{Ran}\hskip 1.9919pt\pi_{[0,c\hskip 0.28436pt\overline{K}_{\varepsilon}]}(\mathcal{L}^{\varepsilon})=1,$
    with $\pi_{[0,c\hskip 0.28436pt\overline{K}_{\varepsilon}]}$ the spectral projection of $\mathcal{L}^{\varepsilon}$ on $[0,c\hskip 0.28436pt\overline{K}_{\varepsilon}]$.
In particular, this result shows that the first eigenpair $(\lambda_{0}^{\varepsilon},u_{0}^{\varepsilon})$ is well defined, and also provides the upper bound $0\leqslant\lambda_{0}^{\varepsilon}\leqslant c\overline{K}_{\varepsilon}$ in terms of the quantity introduced in (1.10).
The second result describes the exit event. In this work, we consider that the domain is actually a small modification of the disk (the unit ball in dimension 2), denoted by $\widetilde{\Omega}_{\varepsilon}$, see Figure 2 for a schematic representation. As will become clear below, the motivation for introducing this specific domain is to obtain a precise asymptotic expansion of the first eigenvalue and to determine the law of the first exit point.
Heuristic construction of approximate solutions to (1.5).
In order to define the modified domain $\widetilde{\Omega}_{\varepsilon}$, we need to introduce the quasimode which is used to approximate the first eigenfunction $u_{0}^{\varepsilon}$. The discussion below is informal: rigorous justifications of the quality of the quasimode are given in the forthcoming sections.
Let us first recall that, for a self-adjoint operator $T$ on a Hilbert space $\mathcal{H}$, an $\mathcal{H}$-normalized element $u$ such that $\|\pi_{[b,+\infty)}(T)u\|$ is small is called a quasi-mode for the spectrum in $[0,b]$ of $T$. Bounds on the spectral projection $\|\pi_{[b,+\infty)}(T)u\|$ are typically obtained from bounds on the quadratic form associated with $T$, see [23, Lemma 2.4.5] (as well as Lemma 3.4 below in the context of this work).
The use of quasimodes to analyze the asymptotic spectral properties of operator is standard in semiclassical analysis (see [10, 21, 13]). We adopt here a similar strategy, but in a different context since the parameter $\varepsilon$ is not a semiclassical parameter. Recall that the eigenvalue problem of interest is given by (1.5), with $|\Gamma^{\varepsilon}_{\mathbf{D}}|\rightarrow 0$ (in view of (1.3)). The operator $\mathcal{L}^{\varepsilon}$ defined in Theorem 1.1 can therefore be thought of as a perturbation of the Laplacian with Neumann boundary conditions. In particular, one expects that $\lambda_{0}^{\varepsilon}\rightarrow 0$ when $\varepsilon\rightarrow 0$ (which is compatible with the intuition that the mean exit time should go to infinity in this regime, see (1.7)). The $L^{2}$ normalized eigenvector associated with the first eigenvalue 0 of the Neumann Laplacian is (up to a sign) the constant function $-1/\sqrt{\pi}$. This suggests that $u_{0}^{\varepsilon}$ should be a perturbation of $-1/\sqrt{\pi}$, at least sufficiently far away from the Dirichlet zones. This motivates introducing the function $\overline{u}^{\varepsilon}=u_{0}^{\varepsilon}+1/\sqrt{\pi}$. In view of (1.5), the function $\overline{u}^{\varepsilon}$ should satisfy, at leading order in $\varepsilon$ and away from the Dirichlet regions,
(1.15) $\left\{\begin{array}[]{rll}\Delta\overline{u}^{\varepsilon}&\!\!\!\!\displaystyle=\frac{\lambda_{0}^{\varepsilon}}{\sqrt{\pi}}&\quad\text{ in }\Omega,\\ \partial_{n}\overline{u}^{\varepsilon}&\!\!\!\!=0&\quad\text{ on }\Gamma^{\varepsilon}_{\mathbf{N}}.\end{array}\right.$
This suggests to rewrite $\overline{u}^{\varepsilon}$ as
$\overline{u}^{\varepsilon}=-\frac{\lambda_{0}^{\varepsilon}}{\sqrt{\pi}}f$
for some function $f$ to be determined. For a single exit region centered on $x^{(1)}\in\partial\Omega$, the function $f$ formally satisfies (by sending $\varepsilon$ to 0 in (1.15))
(1.16) $\left\{\begin{array}[]{rll}\Delta f&\!\!\!\!=-1&\quad\text{ in }\Omega,\\ \partial_{n}f&\!\!\!\!=-|\Omega|\delta_{x^{(1)}}&\quad\text{ on }\partial\Omega,\end{array}\right.$
where the Dirac mass on the boundary comes from the compatibility condition
$\int_{\partial\Omega}\partial_{n}f=\int_{\Omega}\Delta f=-|\Omega|.$
In view of [1, Equation (3.13)], the solution $f$ to (1.16) is
$f(x)=\log\left|x-x^{(1)}\right|+\frac{1-|x|^{2}}{4},$
where ||$\cdot$|| ⋅ | denotes the Euclidean norm. For $N=1$, an approximation of the solution $u_{0}^{\varepsilon}$ to (1.5)–(1.6) is therefore
$\varphi^{\varepsilon}=-\frac{1}{\sqrt{\pi}}-\frac{\lambda_{0}^{\varepsilon}}{\sqrt{\pi}}f.$
The value of $\lambda_{0}^{\varepsilon}$ allowing to satisfy the Dirichlet boundary conditions at a distance $\mathrm{e}^{-1/K_{\varepsilon}^{(1)}}$ of $x^{(1)}$ is approximately determined by the condition $01+\lambda_{0}^{\varepsilon}\log|x-x^{(1)}|=$ 0, so that $\lambda_{0}^{\varepsilon}=K_{\varepsilon}^{(1)}$ (where we recall that the parameters $(K^{(k)}_{\varepsilon})_{k=1,\ldots,N}$ are related to the size of the absorbing windows, see (1.3)).
When there are $N$ absorbing windows, the above analysis suggests that $u_{0}^{\varepsilon}$ solution to (1.5)–(1.6) should be well approximated away from the Dirichlet regions by
(1.17) ${\varphi}^{\varepsilon}:=-\frac{1}{\sqrt{\pi}}-\frac{1}{\sqrt{\pi}}\sum_{k=1}^{N}K^{(k)}_{\varepsilon}f_{k},\qquad f_{k}(x)=\log\left|x-x^{(k)}\right|+\frac{1-|x|^{2}}{4}.$
The function $\varphi^{\varepsilon}$ converges pointwise to $-1/\sqrt{\pi}$ in $\Omega$ in the limit $\varepsilon\to 0$, and, for a fixed small $\varepsilon$, it is negative in a large central part of $\Omega$, while it diverges to +$\infty$+ ∞ in the vicinity of the boundary points $x^{(k)}$. Note that different weights are introduced in front of each of the functions $f_{k}$. Heuristically, the idea is that $\varphi^{\varepsilon}(x)=0$ approximately translates into $K_{\varepsilon}^{(k)}f_{k}(x)=1$ around the exit $x^{(k)}$ (since only the function $f_{k}$ diverges, the other ones remaining bounded as $x$ approaches the boundary). Since $(1-|x|^{2})/4$ is bounded and $K^{(k)}_{\varepsilon}\rightarrow 0$ as $\varepsilon\rightarrow 0$, the condition $K_{\varepsilon}^{(k)}f_{k}(x)=1$ approximately means that $\left|x-x^{(k)}\right|$ is of order $\mathrm{e}^{-1/K_{\varepsilon}^{(k)}}$ for $x\in\Gamma^{\varepsilon}_{\mathbf{D}_{k}}$ (see Lemma 1.2 below for precise statements). This is indeed consistent with (1.3). The computation of $\Delta\varphi^{\varepsilon}=\overline{K}_{\varepsilon}/\sqrt{\pi}$ suggests that the eigenvalue $\lambda_{0}^{\varepsilon}$ is $\overline{K}_{\varepsilon}$ defined in (1.10) at leading order in $\varepsilon$.
Construction of the modified domain.
We are now in position to precisely introduce the modified domain $\widetilde{\Omega}_{\varepsilon}$, and the associated modification of the eigenvalue problem (1.5) that we will consider in this work. The modified domain is defined by:
(1.18) $\widetilde{\Omega}_{\varepsilon}=\Omega\cap\left(\varphi^{\varepsilon}\right)^{-1}(-\infty,0].$
The associated $N$ Dirichlet regions around the exit points $x^{(k)}$ are denoted by $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}$ for $k\in\{1,\dots,N\}$, with $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}$ defined as the connected component of $\Omega\cap\left(\varphi^{\varepsilon}\right)^{-1}\{0\}$ which is the closest to $x^{(k)}$. We refer to Figure 2 for a schematic representation in the case $N=3$. We are thus interested in the exit event $(\widetilde{\tau}_{\varepsilon},X_{\widetilde{\tau}_{\varepsilon}})$, where
(1.19) $\widetilde{\tau}_{\varepsilon}:=\inf\left\{t\geqslant 0,\hskip 2.84544ptX_{t}\notin\widetilde{\Omega}_{\varepsilon}\right\}.$
$x^{(2)}$$x^{(3)}$$x^{(1)}$$\Gamma^{\varepsilon}_{\mathbf{D}_{2}}$$\Gamma^{\varepsilon}_{\mathbf{D}_{1}}$$\Gamma^{\varepsilon}_{\mathbf{D}_{3}}$$\Gamma^{\varepsilon}_{\mathbf{N}}$$\Gamma^{\varepsilon}_{\mathbf{N}}$$\Gamma^{\varepsilon}_{\mathbf{N}}$$\widetilde{\Omega}_{\varepsilon}$
Figure 2. The domain $\widetilde{\Omega}_{\varepsilon}$ for $N=3$
As quantified in Lemmas 1.2 below, the domain $\widetilde{\Omega}_{\varepsilon}$ and the associated exit regions are small modifications of $\Omega$ and of the exit regions (1.3), and it is therefore expected that the first exit time and first exit point distribution are not substantially changed when passing from $\Omega$ to $\widetilde{\Omega}_{\varepsilon}$ (for the sake of conciseness, this is however not formally proven in this article, and left for future work). In particular, as explained below, the results stated on $\mathcal{L}^{\varepsilon}$ in Theorem 1.1 also hold for ~$\widetilde{\mathcal{L}}^{\varepsilon}$, which is defined similarly as $\mathcal{L}^{\varepsilon}$ but on the modified domain $\widetilde{\Omega}_{\varepsilon}$.
The following lemma follows directly from Lemmas 2.3 and 2.4 (with $\alpha=1$), and shows that the exit regions on the modified domain $\widetilde{\Omega}_{\varepsilon}$ are indeed asymptotically close to the exit regions (1.3) introduced on the original domain $\Omega$.
Lemma 1.2.
There exists $C_{-}>C_{+}>0$ such that, for any ${\varepsilon}\in(0,\varepsilon_{0})$,
$\left(\bigcup\limits_{k=1}^{N}\mathbf{B}\left(x^{(k)},r^{(k)}_{\varepsilon,-}\right)\right)\cap\Omega\subset({\varphi}^{\varepsilon})^{-1}\left([0,+\infty[\right)\cap\Omega\subset\left(\bigcup\limits_{k=1}^{N}\mathbf{B}\left(x^{(k)},r^{(k)}_{\varepsilon,+}\right)\right)\cap\Omega,$
where ${r^{(k)}_{\varepsilon,-}}=\mathrm{e}^{-C_{-}/K^{(k)}_{\varepsilon}}$ and ${r^{(k)}_{\varepsilon,+}}=\mathrm{e}^{-{C}_{+}/K^{(k)}_{\varepsilon}}$ (see (2.13) below).
According to Lemma 1.2, the function ${\varphi}^{\varepsilon}$ vanishes in $\Omega$ on $N$ disjoint curves $\left(\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}\right)_{k=1\ldots,N}$, contained in an annulus centered at $x^{(k)}$, with radii $r^{(k)}_{\varepsilon,-}$ and $r^{(k)}_{\varepsilon,+}$. More precisely, for $k\in\{1,\dots,N\}$, the curve $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}\subset\left(\varphi^{\varepsilon}\right)^{-1}\{0\}\cup\partial\Omega$ is the connected component of $\Omega\cap(\varphi^{\varepsilon})^{-1}\{0\}$ closest to $x^{(k)}$; see Lemma 3.4 (which requires further reducing $\varepsilon_{0}$ to ensure that (3.2) holds). The boundary of $\widetilde{\Omega}_{\varepsilon}$ is thus composed of two parts:
(1.20) $\partial\widetilde{\Omega}_{\varepsilon}=\overline{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}}\cup\overline{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}},\qquad\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}=\bigcup_{k=1}^{N}\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}},\qquad\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}\cap\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}=\emptyset,$
where $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}\subset\partial\Omega$ is the reflecting part.
Our goal now is to examine the narrow escape problem on the domain $\widetilde{\Omega}_{\varepsilon}$; more precisely to study the following eigenvalue problem, which is a modification of (1.5) where $\Omega$ is replaced by $\widetilde{\Omega}_{\varepsilon}$:
(1.24) $\left\{\begin{array}[]{ll}-\Delta\widetilde{u}_{0}^{\varepsilon}&=\widetilde{\lambda}_{0}^{\varepsilon}\,\widetilde{u}_{0}^{\varepsilon}\,\,\text{in}\,\,\widetilde{\Omega}_{\varepsilon},\\ \partial_{n}\widetilde{u}_{0}^{\varepsilon}&=0\,\,\,\text{on}\,\,\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}},\\ \hskip 14.22636pt\widetilde{u}_{0}^{\varepsilon}&=0\,\,\,\text{on}\,\,\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}.\end{array}\right.$
The eigenfunction is normalized so that
(1.25) $\int_{\widetilde{\Omega}_{\varepsilon}}\left(\widetilde{u}_{0}^{\varepsilon}(x)\right)^{2}\,\mathrm{d}x=1,\qquad\widetilde{u}_{0}^{\varepsilon}<0\textrm{ on}\widetilde{\Omega}_{\varepsilon}.$
In the following, we denote by ~$\widetilde{\mathcal{L}}^{\varepsilon}$ the operator associated with (1.24), namely minus the Laplacian with the mixed boundary conditions on the modified domain $\widetilde{\Omega}_{\varepsilon}$. The unique quasi-stationary distribution of the process (1.1) in $\widetilde{\Omega}_{\varepsilon}$ is
$\widetilde{\nu}_{0}^{\varepsilon}(\mathrm{d}x)=\frac{\widetilde{u}_{0}^{\varepsilon}(x)\hskip 2.84544pt\mathrm{d}x}{\displaystyle\int_{\Omega}\widetilde{u}_{0}^{\varepsilon}(y)\hskip 2.84544pt\mathrm{d}y}.$
Working on the modified domain $\widetilde{\Omega}_{\varepsilon}$ allows us to obtain a precise asymptotic expansion of the first eigenvalue and eigenvector $(\widetilde{\lambda}_{0}^{\varepsilon},\widetilde{u}_{0}^{\varepsilon})$, using the fact that the quasimode $\varphi^{\varepsilon}$ is by construction in the domain of ~$\widetilde{\mathcal{L}}^{\varepsilon}$; in particular, is satisfies the Neumann (resp. Dirichlet) boundary conditions on $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}$ (resp. $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}$), see (2.14) below for the Neumann boundary condition. Moreover,
(1.26) for any $k\in\{1,\ldots,N\}$, the curves $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}$ and $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}$ meet at an angle equal to $\frac{\pi}{2}$
since $\partial_{n}\varphi^{\varepsilon}$ vanishes at the intersection points between $\partial\Omega$ (whose normal is $\vec{n}$) and $\left(\varphi^{\varepsilon}\right)^{-1}\{0\}$ (whose normal is $\nabla\varphi^{\varepsilon}$). This implies that the normal derivative of the first eigenvector $\widetilde{u}^{\varepsilon}_{0}$ is in $L^{2}(\partial\widetilde{\Omega}_{\varepsilon})$ (see the proof of Lemma 4.2 and [33, Proposition 32] for more details), and this will be useful in particular to prove that $\partial_{n}\varphi^{\varepsilon}$ is an excellent approximation of $\partial_{n}\widetilde{u}_{0}^{\varepsilon}$ on $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}$.
Since $-\Delta f_{k}=1$ away from $x^{(k)}$ (see (2.15) below), one immediately gets that the following equality holds pointwise in $\widetilde{\Omega}_{\varepsilon}$ (using the notation $\overline{K}_{\varepsilon}$ introduced in (1.10)):
$-\Delta{\varphi}^{\varepsilon}=-\frac{1}{\sqrt{\pi}}\sum_{k=1}^{N}K^{(k)}_{\varepsilon}=\overline{K}_{\varepsilon}{\varphi}^{\varepsilon}+\mathrm{o}(1).$
Note that we crucially use here that $\varphi^{\varepsilon}$ is bounded on $\widetilde{\Omega}_{\varepsilon}$ uniformy in $\varepsilon\in(0,\varepsilon_{0})$. This calculation suggests that $\widetilde{\lambda}_{0}^{\varepsilon}$ is at dominant order equal to $\overline{K}_{\varepsilon}$, as made precise in Theorem 1.3. Let us recall that this result characterizes the law of the first exit time $\widetilde{\tau}_{\varepsilon}$, which is indeed exponential with parameter $\widetilde{\lambda}_{0}^{\varepsilon}$.
Theorem 1.3.
The first eigenvalue $\widetilde{\lambda}_{0}^{\varepsilon}$ of the system (1.24) satisfies, in the limit $\varepsilon\to 0$,
(1.27) $\widetilde{\lambda}_{0}^{\varepsilon}=\overline{K}_{\varepsilon}+\mathrm{O}\left(\overline{K}^{2}_{\varepsilon}\right).$
The final result concerns the law of the first exit point $X_{\widetilde{\tau}_{\varepsilon}}$, which is given by (1.8) with $u_{0}^{\varepsilon}$ replaced by $\widetilde{u}_{0}^{\varepsilon}$, and $\Omega$ replaced by $\widetilde{\Omega}_{\varepsilon}$. Indeed, one can prove the following asymptotic behavior of the normal derivative of $\widetilde{u}_{0}^{\varepsilon}$ as $\varepsilon\to 0$.
Theorem 1.4.
Let $k\in\{1,\dots,N\}$. Let $\widetilde{u}_{0}^{\varepsilon}$ be an eigenfunction of ~$\widetilde{\mathcal{L}}^{\varepsilon}$ associated with the first eigenvalue $\widetilde{\lambda}_{0}^{\varepsilon}$ (see (1.24)) and normalized as (1.25). Then, in the limit $\varepsilon\rightarrow 0$,
(1.28) $\int_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}}\nabla\widetilde{u}_{0}^{\varepsilon}\cdot\vec{n}=\sqrt{\pi}\,\overline{K}_{\varepsilon}+\mathrm{O}\left(\overline{K}_{\varepsilon}^{2}\right).$
Moreover, in the limit $\varepsilon\rightarrow 0$, for all $k\in\{1,\dots,N\}$,
(1.29) $\int_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}}\nabla\widetilde{u}_{0}^{\varepsilon}\cdot\vec{n}=\sqrt{\pi}{K}^{(k)}_{\varepsilon}+\mathrm{O}\left(\overline{K}^{3/2}_{\varepsilon}\right).$
As a consequence, in the limit $\varepsilon\rightarrow 0$, for all $k\in\{1,\dots,N\}$,
(1.30) $\mathbb{P}_{\widetilde{\nu}_{0}^{\varepsilon}}\left[X_{\widetilde{\tau}_{\varepsilon}}\in\Gamma^{\varepsilon}_{\mathbf{D}_{k}}\right]=\frac{{K}^{(k)}_{\varepsilon}}{\overline{K}_{\varepsilon}}+\mathrm{O}\left(\sqrt{\overline{K}_{\varepsilon}}\right).$
Note that the expression (1.28) is not obtained by summing (1.29) over $k\in\{1,\ldots,N\}$, as the error term in (1.28) is smaller than the one in (1.29).
In order for this result to be useful, the remainder term of order $\overline{K}_{\varepsilon}^{1/2}$ in (1.30) should be small compared to $K^{(k)}_{\varepsilon}/\overline{K}_{\varepsilon}$, that is ${K}^{(k)}_{\varepsilon}$ is of order at most $\overline{K}_{\varepsilon}^{3/2}$. This can be illustrated on two prototypical situations:
  • in a case where all the windows shrinks with the same scaling, a possible choice in (1.3) is $\mathrm{e}^{-1/K^{(k)}_{\varepsilon}}=a_{k}\varepsilon$ for some positive $a_{k}$, i.e. $K^{(k)}_{\varepsilon}=-\frac{1}{\log a_{k}+\log\varepsilon}$: in this case, $\mathbb{P}_{\widetilde{\nu}_{0}^{\varepsilon}}\left[X_{\widetilde{\tau}_{\varepsilon}}\in\Gamma^{\varepsilon}_{\mathbf{D}_{k}}\right]$ converges to $1/N$ whatever the values of $a_{k}$;
  • let us next consider a case where the windows shrink at different scales, with $\mathrm{e}^{-1/K^{(k)}_{\varepsilon}}=\varepsilon^{a_{k}}$ for some positive $a_{k}$, i.e. $K^{(k)}_{\varepsilon}=-\frac{1}{a_{k}\log\varepsilon}$: in this case, $\mathbb{P}_{\widetilde{\nu}_{0}^{\varepsilon}}\left[X_{\widetilde{\tau}_{\varepsilon}}\in\Gamma^{\varepsilon}_{\mathbf{D}_{k}}\right]$ converges to $\frac{1/a_{k}}{\sum_{k=1}^{N}1/a_{k}}$; windows which shrink at a slower scale are more likely exits.
1.3. Bibliographic comments and perspectives
As already mentioned in the introduction, the narrow escape problem attracted a lot of attention from the physics community, and early works on the asymptotic behavior of the mean first exit time can for example be found in [19, 49, 47, 48], and in [5, 18] where the authors are particularly interested in studying the role of the initial condition. These works typically rely on formal expansions of Neumann Green’s function for the Laplacian in the domain of interest, with Dirac masses at the escape points, and on results from potential theory [26, 2]. In particular, explicit results can be obtained for domains with specific forms for which these Green’s functions are analytically known (disks, balls, squares, etc). These early results have been extended to general domains and a finite number of infinitely small absorbing windows in [42, 8], using the method of matched asymptotic expansions. Similar techniques are used in [34, 35] to study another geometry, namely a domain which is composed of a relatively big head and several absorbing narrow necks, which leads to Neumann–Robin instead of a Dirichlet boundary condition on the small exit regions. Rigorous derivations of the first and second order terms in the asymptotic expansion of the mean first exit time have been obtained in [1] using layer potential techniques from [2]. More recently, mathematical proofs of the results from [8] have been obtained in [7], where the authors prove the asymptotic expansion rigorously using subsolutions and supersolutions.
As mentioned earlier, our results differ from what has been done previously in two ways. First, our work is, to the best of our knowledge, the first to provide results on the first exit point distribution. Notice that this requires very precise asymptotic estimates, since this exit point distribution is expressed in terms of the normal derivative of the first eigenvector on the boundary, see (1.8). Moreover, the mathematical approach and tools used in our work to study the narrow escape problem (namely the quasi-stationary distribution approach to metastability and spectral techniques in the spirit of those used in semi-classical analysis) also seem to be new. As mentioned above, semi-classical techniques have been used a lot to study the exit problem in the case of an energetic barrier in the small temperature regime, see for example [22, 23, 31, 15, 12, 28, 29, 30, 38, 33, 39]: our work is a first attempt to adapt these techniques to the case of entropic barriers. A major difference between the energetic and the entropic case is that for entropic barriers, the 1-eigenforms do not concentrate around the exit region as fast as for energetic barriers (there are no counterparts to Agmon estimates). We circumvent this difficulty by using an excellent quasimode of the 0-eigenform, as outlined above.
This work opens many perspectives that we would like to explore in the future. First, it would be interesting to prove that modifying the domain from $\Omega$ to $\tilde{\Omega}_{\varepsilon}$ (see the previous section for details) indeed does not modify the asymptotic results that we have obtained on the exit event. Moreover, it should be possible to apply the same method to study the narrow escape problem in a general bounded domain in $\mathbb{R}^{d}$. Besides, it would be interesting to investigate if similar techniques can be used to study the narrow escape problem for the underdamped Langevin dynamics (commonly used in molecular dynamics): this is more challenging since the associated infinitesimal generator is non-selfadjoint and hypoelliptic, and this has major implications on the functional setting (in particular to state the boundary conditions, see e.g. [32] for more details) and on the spectral properties of the operator. Finally, as already mentioned above, the quasi-stationary approach to metastability is deeply linked with a class of numerical methods to efficienlty sample the exit event of a metastable dynamics [11, 41]. For example, it would be interesting to study how the results we have obtained can be used to obtain generalizations of the temperature accelerated dynamics algorithm [50] in our context (the size of the exit windows playing the role of the temperature).
1.4. Outline of the paper
In Section 2, we investigate the spectral properties of the Laplacian with mixed Dirichlet–Neumann boundary conditions. In Section 3, we present the results regarding the law of the first exit time. Section 4 focuses on the law of the first exit point. In Appendix A, we present the spectral analysis of the Laplacian on $p$-forms with mixed tangential-normal boundary conditions.
Some of the theoretical results on the spectral properties of the $p$-Laplacian with mixed tangential-normal boundary conditions are illustrated along the manuscript by numerical experiments, using in particular Raviart–Thomas orthogonal finite elements (a.k.a. Nedelec finite elements) in order to obtain stable discretizations [4].
2. Spectral analysis
In this section, we perform the spectral analysis of the operator $\mathcal{L}^{\varepsilon}$. We define its domain in Section 2.1. The proof of Theorem 1.1 is given in Section 2.2. Section 2.3 provides a result concerning the second eigenvalue of the operator $\mathcal{L}^{\varepsilon}$. A numerical illustration of our theoretical findings is given in Section 2.4.
In all this section, we consider the operator $\mathcal{L}^{\varepsilon}$ with domain $\mathcal{D}(\mathcal{L}^{\varepsilon})$, see (1.12). Our aim is (i) to introduce the quasimode $\varphi^{\epsilon}$ which will then be used to construct the modified domain $\widetilde{\Omega}_{\varepsilon}$, and (ii) to provide spectral properties of the operator $\mathcal{L}^{\varepsilon}$, that will actually also hold for the mixed Laplacian on the modified domain $\widetilde{\Omega}_{\varepsilon}$, denoted by ~$\widetilde{\mathcal{L}}^{\varepsilon}$ in the following.
2.1. Definition and first properties of the operator $\mathcal{L}^{\varepsilon}$
We begin by studying the spectral properties of the operator $\mathcal{L}^{\varepsilon}$, namely the mixed Laplacian in the unit disk $\Omega$ (see Figure 1). Recall that the boundary of the domain $\Omega$ contains $N$ small disjoint exit regions $(\Gamma^{\varepsilon}_{\mathbf{D}_{k}})_{k=1,\dots,N}$ of respective sizes $\mathrm{e}^{-1/K^{(k)}_{\varepsilon}}$ (recall (1.3)). The boundary $\partial\Omega$ is thus composed of two parts: $\Gamma^{\varepsilon}_{\mathbf{D}}={\cup}_{k=1}^{N}\Gamma^{\varepsilon}_{\mathbf{D}_{k}}$ and $\Gamma^{\varepsilon}_{\mathbf{N}}$ which is the reflecting part. We consider the mixed Dirichlet–Neumann Laplacian eigenvalue problem on $\Omega$, formulated as:
(2.4) $\left\{\begin{array}[]{ll}\mathcal{L}^{\varepsilon}u^{\varepsilon}&=\lambda^{\varepsilon}u^{\varepsilon}\,\,\text{in}\,\,\Omega,\\ \partial_{n}u^{\varepsilon}&=0\,\,\,\text{on}\,\,\Gamma^{\varepsilon}_{\mathbf{N}},\\ \hskip 14.22636ptu^{\varepsilon}&=0\,\,\,\text{on}\,\,\Gamma^{\varepsilon}_{\mathbf{D}_{k}},\qquad\forall k\in\{1,\dots,N\}.\end{array}\right.$
Let us introduce
(2.5) $H^{1}_{0,\Gamma^{\varepsilon}_{\mathbf{D}}}(\Omega):=\left\{u\in H^{1}(\Omega),\,\,u_{|_{\Gamma^{\varepsilon}_{\mathbf{D}}}}=0\right\},$
the space of functions in $H^{1}(\Omega)$ whose trace vanishes on $\Gamma^{\varepsilon}_{\mathbf{D}}$.
Proposition 2.1.
The quadratic form
(2.6) $Q(u)=\int_{\Omega}{|\nabla u|}^{2}$
with form domain $\mathcal{D}(Q)=H^{1}_{0,\Gamma^{\varepsilon}_{\mathbf{D}}}(\Omega)$ is closed and symmetric. Its Friedrichs extension is the operator $\mathcal{L}^{\varepsilon}=-\Delta$ with domain
(2.7) $\mathcal{D}(\mathcal{L}^{\varepsilon})=\left\{u\in H^{1}(\Omega),\,\,\Delta u\in L^{2}(\Omega),\,\,\partial_{n}u_{|_{\Gamma^{\varepsilon}_{\mathbf{N}}}}=0,\,\,u_{|_{\Gamma^{\varepsilon}_{\mathbf{D}}}}=0\right\}.$
The operator $(\mathcal{L}^{\varepsilon},\mathcal{D}(\mathcal{L}^{\varepsilon}))$ is nonnegative and selfadjoint.
Proof.
Let us consider the following quadratic form
$u\in C_{\Gamma^{\varepsilon}_{\mathbf{D}}}^{\infty}(\Omega)\mapsto\int_{\Omega}{|\nabla u|}^{2},$
where $C_{\Gamma^{\varepsilon}_{\mathbf{D}}}^{\infty}(\Omega)$ is the set of functions belonging to $C^{\infty}(\overline{\Omega})$ vanishing on $\Gamma^{\varepsilon}_{\mathbf{D}}$. It is symmetric, nonnegative and closable. Its closure for the $H^{1}$ norm is the quadratic form
$Q:u\in H^{1}_{0,\Gamma^{\varepsilon}_{\mathbf{D}}}(\Omega)\mapsto\int_{\Omega}{|\nabla u|}^{2}.$
Let $\mathcal{L}^{\varepsilon}$ be the Friedrichs extension associated with the quadratic form $Q$. This operator is nonnegative self-adjoint on $L^{2}(\Omega)$, and defined on the domain
$\mathcal{D}(\mathcal{L}^{\varepsilon})=\left\{u\in H^{1}_{0,\Gamma^{\varepsilon}_{\mathbf{D}}}(\Omega)\,\middle|\,\exists f\in L^{2}(\Omega),\,\forall v\in H^{1}_{0,\Gamma^{\varepsilon}_{\mathbf{D}}},\,Q(u,v)=\left\langle f,v\right\rangle_{L^{2}(\Omega)}\right\}$
by
$-\Delta u=f.$
It is then standard to check that, equivalently, the domain of the Laplacian with mixed Dirichlet–Neumann boundary conditions is given by (2.7). Indeed, one first shows that $-\Delta u=f$ in the sense of distributions, hence $\Delta u\in L^{2}(\Omega)$ since $f\in L^{2}(\Omega)$; and next sees that $\partial_{n}u=0$ on $\Gamma^{\varepsilon}_{\mathbf{N}}$ as elements of $H^{-1/2}(\partial\Omega)$ by an integration by parts. ∎
Remark 2.2.
Concerning the regularity of the functions $u\in\mathcal{D}(\mathcal{L}^{\varepsilon})$, let us recall that because of the mixed boundary conditions, $u$ is not in $H^{2}(\Omega)$ (as would be the case for pure Dirichlet or Neumann boundary conditions) but only in $H^{3/2-\alpha}(\Omega)$ for any $\alpha>0$ (this result holds for any $C^{1,1}$ domain). Actually it can be shown that $u\in B_{2,\infty}^{3/2}(\Omega)$, where $B_{2,\infty}^{3/2}(\Omega)$ is defined by interpolation: $B_{2,\infty}^{3/2}(\Omega)=[H^{1}(\Omega),H^{2}(\Omega)]_{1/2,\infty}$. We refer for example to [45] for more details.
We are now in position to prove item (i) of Theorem 1.1.
Proof of item (i) of Theorem 1.1.
We already stated in Proposition 2.1 that $\mathcal{L}^{\varepsilon}$ is nonnegative and self-adjoint. It therefore only remains to prove that it has a compact resolvent. This is a consequence of the continuity of the inclusion $\mathcal{D}(\mathcal{L}^{\varepsilon})\subset H^{1}(\Omega)$, and of the compactness of the embedding $H^{1}(\Omega)\subset L^{2}(\Omega)$ (guaranteed by the boundedness of $\Omega$). ∎
A corollary of the above result is that the eigenvalues of the mixed Laplacian $\mathcal{L}^{\varepsilon}$, when ranked in increasing order and counted with their multiplicities, are given by the min-max principle: for $n=0$,
(2.8) $\lambda^{\varepsilon}_{0}={\underset{u\in H^{1}_{0,\Gamma^{\varepsilon}_{\mathbf{D}}}(\Omega)\backslash\{0\}}{\text{inf}}}\left\{\frac{Q(u)}{\|u\|^{2}_{L^{2}(\Omega)}}\right\},$
and, for $n\geqslant 1$,
(2.9) $\lambda^{\varepsilon}_{n}=\underset{{E}_{n}:\,\,\text{dim}\,{E}_{n}=n}{\text{sup}}\,\,{\underset{u\in\left(H^{1}_{0,\Gamma^{\varepsilon}_{\mathbf{D}}}(\Omega)\backslash\{0\}\right)\cap{E}^{\perp}_{n}}{\text{inf}}}\left\{\frac{Q(u)}{\|u\|^{2}_{L^{2}(\Omega)}}\right\}.$
2.2. Proof of item (ii) of Theorem 1.1
We provide in this section the proof of item (ii) of Theorem 1.1, which relies on estimates on the Dirichlet form evaluated at the quasi-mode $\varphi^{\varepsilon}$. The proof is based on various technical estimates, which we state and prove before concluding the section with the actual proof of Theorem 1.1.
The first two lemmas allow us to localize the zero set of the quasi-mode $\varphi^{\varepsilon}$ defined in (1.17). In fact, we write localization results for slightly more general functions
(2.10) $\forall x\in\Omega,\qquad{\varphi}_{\alpha}^{\varepsilon}(x)=-\frac{1}{\sqrt{\pi}}-\frac{1}{\alpha\sqrt{\pi}}\sum_{k=1}^{N}K^{(k)}_{\varepsilon}\left(\text{log}\hskip 2.84544pt|x-x^{(k)}|+\frac{1-|x|^{2}}{4}\right),$
for some scaling parameter $\alpha\in(0,+\infty)$. Note that $\varphi^{\varepsilon}_{1}$ coincides with $\varphi^{\varepsilon}$.
Lemma 2.3.
Fix $\alpha\in(0,+\infty)$ and let ${\varphi}_{\alpha}^{\varepsilon}$ be the function defined in (2.10). Then, for
(2.11) ${C}_{\alpha,-}>\alpha+\dfrac{1}{8}+\frac{\log(2)}{2},$
and for any ${\varepsilon}\in(0,\varepsilon_{0})$, if $|x-x^{(k)}|\leqslant\mathrm{e}^{-C_{\alpha,-}/K^{(k)}_{\varepsilon}}$ for some $k\in\{1,\dots,N\}$, then ${\varphi}_{\alpha}^{\varepsilon}(x)>0$.
Proof.
Consider $x\in\Omega$ such that $|x-x^{(k)}|\leqslant\mathrm{e}^{-C_{\alpha,-}/K^{(k)}_{\varepsilon}}$ for some $k\in\{1,\dots,N\}$ (and for some ${C}_{\alpha,-}>0$ to be determined). Using that $|x-x^{(k^{\prime})}|\leqslant 2$ for all $k^{\prime}\in\{1,\dots,N\}$ with $k^{\prime}\neq k$, we obtain, for any ${\varepsilon}\in(0,\varepsilon_{0})$,
${\varphi}^{\varepsilon}(x)$ $\geqslant-\frac{1}{\sqrt{\pi}}+\frac{{C}_{\alpha,-}}{\alpha\sqrt{\pi}}-\left(\sum_{\begin{subarray}{c}k^{\prime}=1\\ k^{\prime}\neq k\end{subarray}}^{N}K^{(k^{\prime})}_{\varepsilon}\right)\frac{\text{log}(2)}{\alpha\sqrt{\pi}}-\left(\sum_{\begin{subarray}{c}k=1\end{subarray}}^{N}K^{(k)}_{\varepsilon}\right)\frac{1}{4\alpha\sqrt{\pi}}$
$\geqslant\frac{1}{{\sqrt{\pi}}}\left(-1+\frac{{C}_{\alpha,-}}{\alpha}-\frac{\log(2)}{\alpha}\overline{K}_{\varepsilon}-\frac{\overline{K}_{\varepsilon}}{4\alpha}\right)$
$\geqslant\frac{1}{{\sqrt{\pi}}}\left(-1+\frac{1}{\alpha}\left[C_{\alpha,-}-\frac{\text{log}(2)}{2}-\frac{1}{8}\right]\right)$
where we used the inequality $\overline{K}_{\varepsilon}<1/2$ implied by (1.11). The right hand side of the last inequality is positive for the choice (2.11). ∎
Lemma 2.4.
Fix $\alpha\in(1/2,+\infty)$ and let ${\varphi}_{\alpha}^{\varepsilon}$ be the function defined in (2.10). Then, for
(2.12) $0<{C}_{\alpha,+}<\alpha-\frac{1}{2},$
and for any ${\varepsilon}\in(0,\varepsilon_{0})$, if $|x-x^{(k)}|\geqslant\mathrm{e}^{-C_{\alpha,+}/K^{(k)}_{\varepsilon}}$ for all $k\in\{1,\dots,N\}$, then ${\varphi}^{\varepsilon}_{\alpha}(x)<0$.
Proof.
Consider $x\in\Omega$ such that $|x-x^{(k)}|\geqslant\mathrm{e}^{-C_{\alpha,+}/K^{(k)}_{\varepsilon}}$ for all $k\in\{1,\dots,N\}$. Introduce $j(x)\in\{1,\ldots,N\}$ such that
$\forall k\in\{1,\dots,N\},\qquad\left|x-x^{(j(x))}\right|\leqslant\left|x-x^{(k)}\right|.$
Using (1.2), it follows $\rho_{0}\leqslant|x^{(k)}-x^{(j(x))}|\leqslant|x^{(k)}-x|+|x-x^{(j(x))}|\leqslant 2|x^{(k)}-x|$, so that $|x-x^{(k)}|\geqslant\rho_{0}/2$ for all $k\neq j(x)$. Then, making use of the inequality $-\log\left(\frac{\rho_{0}}{2}\right)\geqslant 0$, we obtain, for ${\varepsilon}\in(0,\varepsilon_{0})$,
${\varphi}^{\varepsilon}_{\alpha}(x)$ $\leqslant-\frac{1}{\sqrt{\pi}}+\frac{{C}_{\alpha,+}}{\alpha\sqrt{\pi}}-\frac{1}{\alpha\sqrt{\pi}}\sum_{\begin{subarray}{c}k^{\prime}=1\\ k^{\prime}\neq k\end{subarray}}^{N}K_{\varepsilon}^{(k^{\prime})}\log\left(\frac{\rho_{0}}{2}\right)\leqslant\frac{1}{\sqrt{\pi}}\left(-1+\frac{{C}_{\alpha,+}}{\alpha}-\frac{\overline{K}_{\varepsilon}}{\alpha}\log\left(\frac{\rho_{0}}{2}\right)\right).$
The latter quantity is negative in view of (1.11) and (2.12). ∎
In the following, we denote by
(2.13) ${r^{(k)}_{\varepsilon,-}}=\mathrm{e}^{-C_{1,-}/K^{(k)}_{\varepsilon}},\qquad{r^{(k)}_{\varepsilon,+}}=\mathrm{e}^{-C_{1,+}/K^{(k)}_{\varepsilon}},$
the radii which appear respectively in the statements of Lemmas 2.3 and 2.4 for $\alpha=1$. In view of these results, as already stated in the introduction, $\varphi_{\varepsilon}$ vanishes in $\Omega$ on $N$ disjoint curves $(\Gamma^{\varepsilon}_{\mathbf{D}_{k}})_{k=1\ldots,N}$, and for $k\in\{1,\dots,N\}$, the curve $\Gamma^{\varepsilon}_{\mathbf{D}_{k}}$ is located in a neighborhood of $x^{(k)}$ and is contained in an annulus centered at $x^{(k)}$, with radii $r^{(k)}_{\varepsilon,-}$ and $r^{(k)}_{\varepsilon,+}$.
The following results on the functions $f_{k}$ are useful in the proof of item (ii) of Theorem 1.1 below, to get precise estimate on the energy of the quasi-mode $\varphi^{\varepsilon}$
Lemma 2.5.
Let $f_{k}$ be the function defined in (1.17). The function $f_{k}$ satisfies
[…]
and
(2.15) $\Delta f_{k}=-1\text{ in }\Omega.$
Moreover, there exists $C>0$ independent of $k$ such that
(2.16) $\|{f}_{k}\|_{L^{2}(\Omega)}\leqslant C.$
Finally, there exists $C>0$ independent of $k$ and of $\varepsilon$ such that for any ${\varepsilon}\in(0,\varepsilon_{0})$ and $r\in(0,1)$,
(2.17) $\int_{\Omega\setminus{\mathbf{B}}\left(x^{(k)},\,r\right)}{|\nabla f_{k}(x)|}^{2}\,\mathrm{d}x\leqslant C\left(1+\log\left(\frac{2}{r}\right)\right),$
where for any $x\in\Omega$ and $r>0$, $\Omega\setminus{\mathbf{B}}(x,r)$ denotes the complement of the disk ${\mathbf{B}}(x,r)$ in $\Omega$.
Proof.
Let us first prove (2.14). A simple computation shows that
(2.18) $\nabla f_{k}(x)=\frac{x-x^{(k)}}{|x-x^{(k)}|^{2}}-\frac{x}{2}.$
Using that $\vec{n}(x)=\frac{x}{|x|}\,\,\,\text{on}\,\,\partial\Omega$, we obtain that
$\nabla f_{k}(x)\cdot\vec{n}(x)=\frac{x-x^{(k)}}{|x-x^{(k)}|^{2}}\cdot\frac{x}{|x|}-\frac{1}{2}.$
Using polar coordinates, we can write $x^{(k)}=(\text{cos}(\theta_{k}),\text{sin}(\theta_{k}))$ and $x=(\text{cos}(\theta_{x}),\text{sin}(\theta_{x}))$ for $x\in\partial\Omega$ where $\theta_{x},\theta_{k}\in[0,2\pi)$ and $\theta_{x}\neq\theta_{k}$ for $x\in\partial\Omega\setminus\{x^{(k)}\}$. Then,
$\forall x\in\partial\Omega\setminus\{x^{(k)}\},\qquad\nabla f_{k}(x)\cdot\vec{n}(x)=\frac{1-\text{cos}(\theta_{x}-\theta_{k})}{{2(1-\text{cos}(\theta_{x}-\theta_{k}))}}-\frac{1}{2}=0.$
Moreover, since $\frac{x-x^{(k)}}{|x-x^{(k)}|^{2}}=\begin{pmatrix}\partial_{x_{2}}\\ -\partial_{x_{1}}\end{pmatrix}\arctan\left(\frac{x_{2}-x^{(k)}_{2}}{x_{1}-x^{(k)}_{1}}\right)$, one has
${\rm div}\,\left(\frac{x-x^{(k)}}{|x-x^{(k)}|^{2}}\right)=0,$
so that that
$\Delta f_{k}={\rm div}\,(\nabla f_{k})=-1.$
The estimate (2.16) is immediate from the definition of $f_{k}$.
Let us finally prove (2.17). Denote by ${\mathbf{B}}^{\mathrm{c}}(x,r)=\Omega\setminus{\mathbf{B}}(x,r)$ the complement of the disk ${\mathbf{B}}(x,r)$ in $\Omega$. From (2.18), a Cauchy–Schwarz inequality gives
$|\nabla f_{k}|^{2}$ $=\frac{1}{|x-x^{(k)}|^{2}}-\frac{x\cdot(x-x^{(k)})}{|x-x^{(k)}|^{2}}+\frac{|x|^{2}}{4}\leqslant 2\left(\frac{1}{|x-x^{(k)}|^{2}}+\frac{|x|^{2}}{4}\right).$
Then, using that $|x|\leqslant 1$ on ${\mathbf{B}}^{\mathrm{c}}\left(x^{(k)},\,r\right)\subset\Omega$,
$\int_{{\mathbf{B}}^{\mathrm{c}}\left(x^{(k)},r\right)}{|\nabla f_{k}|}^{2}\,\mathrm{d}x$ $\leqslant 2\left(\int_{{\mathbf{B}}^{\mathrm{c}}\left(x^{(k)},\,r\right)}\frac{1}{|x-x^{(k)}|^{2}}\,\mathrm{d}x+\frac{\pi}{4}\right).$
To estimate the first term in the right-hand side, we use polar coordinates centered at $x^{(k)}$ and the fact that ${\mathbf{B}}^{\mathrm{c}}\left(x^{(k)},r\right)\subset\Omega\subset\mathbf{B}\left(x^{(k)},2\right)$ to get
$\int_{{\mathbf{B}}^{\mathrm{c}}\left(x^{(k)},\,r\right)}\frac{1}{|x-x^{(k)}|^{2}}\,\mathrm{d}x\leqslant\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\int_{r}^{2}\frac{1}{r^{\prime}}\,\mathrm{d}r^{\prime}\,\mathrm{d}\theta=\pi\text{log}\left(\frac{2}{r}\right).$
This concludes the proof of (2.17) with $C=2\pi$. ∎
We are now in position to conclude the proof of Theorem 1.1.
Proof of item (ii) of Theorem 1.1.
Let us first prove that
[…]
To this end, we use (2.8) and show that $\mathcal{L}^{\varepsilon}$ admits at least one eigenvalue of order $\mathrm{O}\left(\overline{K}_{\varepsilon}\right)$ by constructing an appropriate test function. Consider
$\frac{1}{2}<\alpha<\frac{7}{8}-\frac{\log(2)}{2},$
so that (2.11) holds with $C_{\alpha,-}=1$ and (2.12) for some positive constant $C_{\alpha,+}$. Recall the definition (2.10) of $\varphi^{\varepsilon}_{\alpha}$, and define $\Omega_{\alpha,\varepsilon}:=\{x\in\Omega,\,{\varphi}^{\varepsilon}_{\alpha}(x)<0\}$ and
(2.20) $\forall x\in\Omega,\qquad{\phi}^{\varepsilon}_{\alpha}(x)={\varphi}^{\varepsilon}_{\alpha}(x){\mathds{1}}_{\Omega_{\alpha,\varepsilon}}(x).$
By construction,
$\Gamma^{\varepsilon}_{\mathbf{D}}=\partial\Omega\cap\left(\bigcup_{k=1}^{N}{\mathbf{B}}\left(x^{(k)},\mathrm{e}^{-1/K_{\varepsilon}^{(k)}}\right)\right)\subset\Omega\setminus\Omega_{\alpha,\varepsilon}.$
so that ${\phi}_{\alpha}^{\varepsilon}$ satisfies Dirichlet boundary conditions on $\Gamma^{\varepsilon}_{\mathbf{D}}$, and therefore ${\phi}_{\alpha}^{\varepsilon}\in H^{1}_{0,\Gamma^{\varepsilon}_{\mathbf{D}}}(\Omega)$. Using (2.9), it follows that
(2.21) $\lambda^{\varepsilon}_{0}\leqslant\frac{Q({\phi}_{\alpha}^{\varepsilon})}{\|{\phi}_{\alpha}^{\varepsilon}\|^{2}_{L^{2}(\Omega)}}.$
An upper bound on $\lambda^{\varepsilon}_{0}$ can thus be obtained from an upper bound on $Q({\phi}^{\varepsilon}_{\alpha})$ and a lower bound on $\|{\phi}^{\varepsilon}_{\alpha}\|^{2}_{L^{2}(\Omega)}$. For the numerator, we use a discrete Cauchy–Schwarz inequality to write
(2.22) $Q({\phi}^{\varepsilon}_{\alpha})$ $\leqslant\frac{N}{\pi\alpha^{2}}\sum_{k=1}^{N}\left(K^{(k)}_{\varepsilon}\right)^{2}\|\nabla f_{k}\|_{L^{2}(\Omega_{\alpha,\varepsilon})}^{2}$
(2.23) $\leqslant\frac{N}{\pi\alpha^{2}}\sum_{k=1}^{N}\left(K^{(k)}_{\varepsilon}\right)^{2}\|\nabla f_{k}\|_{L^{2}(\Omega\setminus\mathbf{B}(x^{(k)},\exp(-1/K_{\varepsilon}^{(k)})))}^{2}\leqslant C\overline{K}_{\varepsilon},$
where the second inequality follows from Lemma 2.3 and the last one from (2.17). For the denominator, we use Lemma 2.4 and (2.16) to write
$\|{\phi}^{\varepsilon}_{\alpha}\|_{L^{2}(\Omega)}=\|{\varphi}^{\varepsilon}_{\alpha}\|_{L^{2}(\Omega_{\alpha,\varepsilon})}$ $\geqslant\frac{|\Omega_{\alpha,\varepsilon}|^{1/2}}{\sqrt{\pi}}-\frac{1}{\alpha\sqrt{\pi}}\sum_{k=1}^{N}K^{(k)}_{\varepsilon}\|f_{k}\|_{L^{2}(\Omega_{\alpha,\varepsilon})}$
(2.24) $\geqslant 1-C\sum_{k=1}^{N}\left(\mathrm{e}^{-C_{\alpha,+}/K_{\varepsilon}^{(k)}}\right)^{2}-C\overline{K}_{\varepsilon},$
since
(2.25) $|\Omega_{\alpha,\varepsilon}|\geqslant\pi\left(1-\frac{1}{2}\sum_{k=1}^{N}\left(\mathrm{e}^{-C_{\alpha,+}/K_{\varepsilon}^{(k)}}\right)^{2}\right).$
For further use, let us notice that, thanks to (2.16), it can easily be shown that $\|{\phi}^{\varepsilon}_{\alpha}\|_{L^{2}(\Omega)}\leqslant 1+C\overline{K}_{\varepsilon}$, so that one actually has, for any $\alpha>1/2$,
(2.26) $\|{\varphi}^{\varepsilon}_{\alpha}\|_{L^{2}(\Omega_{\alpha,\varepsilon})}=1+\mathrm{O}(\overline{K}_{\varepsilon}).$
Gathering (2.22) and (2.24), one obtains that $\lambda_{0}^{\varepsilon}\leqslant c\overline{K}_{\varepsilon}$, for a constant $c$ independent of $\varepsilon$, which yields (2.19).
Let us next prove that $\text{dim}\hskip 1.9919pt\text{Ran}\hskip 0.56917pt\pi_{[0,c\hskip 0.28436pt\overline{K}_{\varepsilon}]}(\mathcal{L}^{\varepsilon})\leqslant 1$. To this end, consider $w\in\mathcal{D}(Q)$ such that $\|w\|_{L^{2}(\Omega)}=1$ and
(2.27) $\int_{\Omega}{|\nabla w(x)|}^{2}\,\mathrm{d}x\leqslant c\hskip 0.28436pt\overline{K}_{\varepsilon},$
and let us prove that the $L^{2}(\Omega)$-distance between $w$ and $\text{Span}(\phi_{\alpha}^{\varepsilon})$ goes to 0 when $\varepsilon\rightarrow 0$, with $\phi^{\varepsilon}_{\alpha}$ defined in (2.20). Denoting by $\overline{\phi}_{\alpha}^{\varepsilon}=\phi_{\alpha}^{\varepsilon}/\|\phi_{\alpha}^{\varepsilon}\|_{L^{2}(\Omega)}$,
$\text{dist}_{L^{2}(\Omega)}\left(w,\text{Span}(\phi_{\alpha}^{\varepsilon})\right)$ $=\left\|w-\left\langle w,\overline{\phi}_{\alpha}^{\varepsilon}\right\rangle_{L^{2}(\Omega)}\overline{\phi}_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega)}$
$\leqslant\|{w}-\left\langle{w}\right\rangle\|_{L^{2}(\Omega)}+\left\|\left\langle{w}\right\rangle-\left\langle{w},{\overline{\phi}}_{\alpha}^{\varepsilon}\right\rangle_{L^{2}(\Omega)}{\overline{\phi}}_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega)},$
where
$\left\langle{w}\right\rangle:=\frac{1}{|\Omega|}\int_{\Omega}w(x)\,\mathrm{d}x=\frac{1}{\pi}\int_{\Omega}w(x)\,\mathrm{d}x.$
The idea behind the previous decomposition is that the leading order term of the function ${{\varphi}}_{\alpha}^{\varepsilon}$ is $-1/\sqrt{\pi}$, which implies that the leading order term of $\left\langle{w},{{\varphi}}_{\alpha}^{\varepsilon}\right\rangle_{L^{2}(\Omega)}{{\varphi}}_{\alpha}^{\varepsilon}$ is $\left\langle{w}\right\rangle$. Concerning the first term, the classical Poincaré–Wirtinger inequality and (2.27) imply that
(2.28) $\|{w}-\left\langle{w}\right\rangle\|_{L^{2}(\Omega)}\leqslant C_{\Omega}\|\nabla{w}\|_{L^{2}(\Omega)}\leqslant C_{\Omega}^{{}^{\prime}}\sqrt{\overline{K}_{\varepsilon}}.$
For the second term, denoting by ${{\varphi}}^{0}=-1/\sqrt{\pi}$, we write
$\left\|\left\langle{w}\right\rangle-\left\langle{w},{\overline{\phi}}_{\alpha}^{\varepsilon}\right\rangle_{L^{2}(\Omega)}{\overline{\phi}}_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega)}$
$=\left\|\left\langle{w}\right\rangle-\left\langle{w},{{\varphi}}^{0}\right\rangle_{L^{2}(\Omega)}{\overline{\phi}}_{\alpha}^{\varepsilon}+\left\langle{w},{{\varphi}}^{0}\right\rangle_{L^{2}(\Omega)}{\overline{\phi}}_{\alpha}^{\varepsilon}-\left\langle{w},{\overline{\phi}}_{\alpha}^{\varepsilon}\right\rangle_{L^{2}(\Omega)}{\overline{\phi}}_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega)}$
$=\left\|\left\langle{w},{{\varphi}}^{0}\right\rangle_{L^{2}(\Omega)}\left({{\varphi}}^{0}-{\overline{\phi}}_{\alpha}^{\varepsilon}\right)+\left\langle{w},{{\varphi}}^{0}-{\overline{\phi}}_{\alpha}^{\varepsilon}\right\rangle_{L^{2}(\Omega)}{\overline{\phi}}_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega)}$
$\leqslant\left\|\left\langle{w},{{\varphi}}^{0}\right\rangle_{L^{2}(\Omega)}\left({{\varphi}}^{0}-{\overline{\phi}}_{\alpha}^{\varepsilon}\right)\right\|_{L^{2}(\Omega)}+\left\|\left\langle{w},{{\varphi}}^{0}-{\overline{\phi}}_{\alpha}^{\varepsilon}\right\rangle_{L^{2}(\Omega)}{\overline{\phi}}_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega)}$
$\leqslant\left|\left\langle{w},{{\varphi}}^{0}\right\rangle_{L^{2}(\Omega)}\right|\left\|{{\varphi}}^{0}-{\overline{\phi}}_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega)}+\|w\|_{L^{2}(\Omega)}\left\|{{\varphi}}^{0}-{\overline{\phi}}_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega)}$
$\leqslant 2\|w\|_{L^{2}(\Omega)}\left\|{{\varphi}}^{0}-{\overline{\phi}}_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega)}\leqslant C\overline{K}_{\varepsilon},$
where we used that $\|{w}\|_{L^{2}(\Omega)}=\|{\overline{\phi}}_{\alpha}^{\varepsilon}\|_{L^{2}(\Omega)}=\|{{\varphi}}^{0}\|_{L^{2}(\Omega)}=1$. For the last inequality, we rely on the bound
$\left\|{{\varphi}}^{0}-{\overline{\phi}}_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega)}$ $\leqslant\left\|{{\varphi}}^{0}-{\overline{\phi}}_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega_{\alpha,\varepsilon})}+\left\|{{\varphi}}^{0}\right\|_{L^{2}(\Omega\setminus\Omega_{\alpha,\varepsilon})}$
$\leqslant\left\|{{\varphi}}^{0}-\varphi_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega_{\alpha,\varepsilon})}+\left\|\varphi_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega_{\alpha,\varepsilon})}-1+\left\|{{\varphi}}^{0}\right\|_{L^{2}(\Omega\setminus\Omega_{\alpha,\varepsilon})}$
$\leqslant\left\|{{\varphi}}^{0}-\varphi_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega_{\alpha,\varepsilon})}+\left\|\varphi_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega_{\alpha,\varepsilon})}-1+\frac{1}{\sqrt{\pi}}\left(\pi-\left|\Omega_{\alpha,\varepsilon}\right|\right),$
where the term $\left\|\varphi_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega_{\alpha,\varepsilon})}-1$ is of order $\overline{K}_{\varepsilon}$ using (2.26), the last term is of order $\overline{K}_{\varepsilon}$ in view of (2.25), and the first one is bounded as
(2.29) $\left\|{{\varphi}}^{0}-{{\varphi}}_{\alpha}^{\varepsilon}\right\|_{L^{2}(\Omega_{\alpha,\varepsilon})}$ $=\frac{1}{\alpha\sqrt{\pi}}\left\|\displaystyle\sum_{k=1}^{N}K^{(k)}_{\varepsilon}f_{k}\right\|_{L^{2}(\Omega_{\alpha,\varepsilon})}\leqslant\frac{1}{\alpha\sqrt{\pi}}\sum_{k=1}^{N}K^{(k)}_{\varepsilon}\left\|f_{k}\right\|_{L^{2}(\Omega)}\leqslant C\overline{K}_{\varepsilon},$
where we used (2.16) for the last inequality. In conclusion, we obtain that, for $\varepsilon\in(0,\varepsilon_{0})$,
$\text{dist}_{L^{2}(\Omega)}\left(w,\text{Span}(\phi_{\alpha}^{\varepsilon})\right)\leqslant C\sqrt{\overline{K}_{\varepsilon}}.$
This implies that $\text{dim}\hskip 1.9919pt\text{Ran}\hskip 0.56917pt\pi_{[0,c\hskip 0.28436pt\bar{K}_{\varepsilon}]}(\mathcal{L}^{\varepsilon})\leqslant 1$, and concludes the proof of item (ii) of Theorem 1.1. ∎
2.3. Estimate on the second eigenvalue of $\mathcal{L}^{\varepsilon}$
We next give a classical estimate on the second eigenvalue of the mixed Dirichlet–Neumann Laplacian eigenvalue problem (1.5), by comparing it to the second eigenvalue of either the Neumann or the Dirichlet Laplacian. This result is used in the proof of Proposition 3.4 (in combination with Lemma 3.1 below). In order to state the result, we introduce the Dirichlet Laplacian eigenvalue problem on $\Omega$:
(2.32) $\left\{\begin{array}[]{rll}-\Delta u^{\mathbf{D}}&\!\!\!\!=\lambda^{\mathbf{D}}u^{\mathbf{D}}&\quad\text{in}\,\,\Omega,\\ u^{\mathbf{D}}&\!\!\!\!=0&\quad\text{on}\,\,\partial\Omega.\end{array}\right.$
Since $\Omega$ is a smooth bounded domain, the spectrum of this operator is discrete and consists of eigenvalues $(\lambda^{\mathbf{D}}_{i})_{i\in\mathbb{N}}$ having finite multiplicities such that
(2.33) $0<\lambda^{\mathbf{D}}_{0}<\lambda^{\mathbf{D}}_{1}\leqslant\lambda^{\mathbf{D}}_{2}\leqslant\lambda^{\mathbf{D}}_{3}\leqslant\lambda^{\mathbf{D}}_{4}\dots,\quad\lambda^{\mathbf{D}}_{k}\xrightarrow[k\to\infty]{}\infty.$
Likewise, let us introduce the Neumann Laplacian eigenvalue problem on $\Omega$:
(2.36) $\left\{\begin{array}[]{rll}-\Delta u^{\mathbf{N}}&\!\!\!\!=\lambda^{\mathbf{N}}u^{\mathbf{N}}&\quad\text{in}\,\,\Omega,\\ \partial_{n}u^{\mathbf{N}}&\!\!\!\!=0&\quad\text{on}\,\,\partial\Omega.\end{array}\right.$
Again, the spectrum of this operator is discrete and consists of eigenvalues $(\lambda^{\mathbf{N}}_{i})_{i\in\mathbb{N}}$ with finite multiplicities, tending to infinity:
$0=\lambda^{\mathbf{N}}_{0}<\lambda^{\mathbf{N}}_{1}\leqslant\lambda^{\mathbf{N}}_{2}\leqslant\lambda^{\mathbf{N}}_{3}\leqslant\lambda^{\mathbf{N}}_{4}\dots,\quad\lambda^{\mathbf{N}}_{k}\xrightarrow[k\to\infty]{}\infty.$
Proposition 2.6.
Let $\lambda^{\varepsilon}_{1}$ be the second eigenvalue of the mixed Dirichlet–Neumann problem (1.5). Then, for any $\varepsilon\in(0,\varepsilon_{0})$,
(2.37) $\lambda^{\mathbf{N}}_{1}\leqslant\lambda^{\varepsilon}_{1}\leqslant\lambda^{\mathbf{D}}_{1},$
where $\lambda^{\mathbf{N}}_{1}$ (resp. $\lambda^{\mathbf{D}}_{1}$) is the second eigenvalue of the Neumann (resp. Dirichlet) problem (see respectively (2.36) and (2.32)).
Proof.
Let us introduce
(2.38) $H^{1}_{0}(\Omega):=\left\{u\in H^{1}(\Omega),\,\,u_{|_{\partial\Omega}}=0\right\},$
the space of functions in $H^{1}(\Omega)$ whose trace vanishes on $\partial\Omega$. The min-max principle implies that
$\lambda^{\mathbf{N}}_{1}$ $=\underset{{\psi_{1}}\in L^{2}(\Omega)\backslash\{0\}}{\text{sup}}\quad\underset{u\in\text{Span}(\psi_{1})^{\perp}}{\underset{u\in H^{1}(\Omega)\backslash\{0\}}{\text{inf}}}\left\{\frac{Q(u)}{\|u\|^{2}_{L^{2}(\Omega)}}\right\}$
$\leqslant\underset{{\psi_{1}}\in L^{2}(\Omega)\backslash\{0\}}{\text{sup}}\quad\underset{u\in\text{Span}(\psi_{1})^{\perp}}{\underset{u\in H^{1}_{0,\Gamma^{\varepsilon}_{\mathbf{D}}}(\Omega)\backslash\{0\}}{\text{inf}}}\left\{\frac{Q(u)}{\|u\|^{2}_{L^{2}(\Omega)}}\right\}=\lambda^{\varepsilon}_{1}$
$\leqslant\underset{{\psi_{1}}\in L^{2}(\Omega)\backslash\{0\}}{\text{sup}}\quad\underset{u\in\text{Span}(\psi_{1})^{\perp}}{\underset{u\in H^{1}_{0}(\Omega)\backslash\{0\}}{\text{inf}}}\left\{\frac{Q(u)}{\|u\|^{2}_{L^{2}(\Omega)}}\right\}=\lambda^{\mathbf{D}}_{1},$
where we used the inclusion $H^{1}_{0}(\Omega)\subset H^{1}_{0,\Gamma^{\varepsilon}_{\mathbf{D}}}(\Omega)\subset H^{1}(\Omega)$. ∎
2.4. Numerical illustration of Theorem 1.1
We numerically study here the eigenvalue problem (1.5) on $\Omega$, in order to illustrate the theoretical results obtained in this section. We choose $K_{\varepsilon}=-1/\ln\varepsilon$, so that the holes have a radius $\varepsilon$. The numerical simulations were performed using FreeFem++ [20]. We used piecewise linear continuous $\mathbb{P}_{1}$ finite elements. The mesh was produced using the automatic mesh generator of FreeFem++, with about $80$ cells to discretize the exit regions $\Gamma_{\mathbf{D}_{k}}$ and $160$ cells to mesh the remaining part of the boundary.
Theorem 1.1 shows that the smallest eigenvalue $\lambda^{\varepsilon}_{0}$ of the problem (1.5) is non-degenerate, and that it tends to 00 as $\varepsilon\rightarrow 0$. Proposition 2.6 states that the second eigenvalue $\lambda_{1}^{\varepsilon}$ is bounded from below (and above) for all $\varepsilon>0$. Moreover, as discussed after (1.6), the first eigenfunction $u^{\varepsilon}_{0}$ does not vanish in $\Omega$ and can therefore be chosen to be negative and with $L^{2}$ norm equal to 1. As can be inferred from the proof of item (ii) of Theorem 1.1, one then expects $u^{\varepsilon}_{0}$ to be close to the quasimode $\varphi^{\varepsilon}$, and thus close to the constant function $-1/\sqrt{\pi}$, as $\varepsilon\rightarrow 0$.
Figure 3. First and second eigenvalues of $\mathcal{L}^{\varepsilon}$ as a function of $\varepsilon$ for $N=1$ and $K_{\varepsilon}=-1/\ln\varepsilon$.
Figure 4. Eigenfunction $u^{\varepsilon}_{0}$ and quasimode ${\varphi}^{\varepsilon}$ along the horizontal cut $\Omega\cap\{y=0\}$ for a single exit point $x^{(1)}=(1,0)$ and $\mathrm{e}^{-1/K_{\varepsilon}^{(k)}}=\varepsilon=0.1$.
[…]
We first consider the case of a single absorbing window centered at $x^{(1)}=(1,0)$. As expected, we observe on Figure 3 that the smallest eigenvalue $\lambda^{\varepsilon}_{0}$ of the problem (1.5) tends to 00 as $\varepsilon\rightarrow 0$, while the second eigenvalue $\lambda_{1}^{\varepsilon}$ is almost constant. In order to compare the first eigenfunction with the quasimode $\varphi^{\varepsilon}$, we represent in Figure 4 the first eigenfunction of the operator $\mathcal{L}^{\varepsilon}$ for $\varepsilon=0.1$ along the line $\Omega\cap\{y=0\}$, together with the associated quasi-mode defined in (1.17). We observe that the two functions are indeed very close, except around the exit region.
To confirm this observation, we also consider the case of two small disjoint absorbing windows $\Gamma^{\varepsilon}_{\mathbf{D}_{1}}$ and $\Gamma^{\varepsilon}_{\mathbf{D}_{2}}$ centered at $x^{(1)}=(1,0)$ and $x^{(2)}=(-1,0)$. We represent in Figure 5 the numerical approximations of the first two eigenfunctions for $\varepsilon=0.1$. We again observe that the first eigenfunction is close to the constant function $-1/\sqrt{\pi}\approx-0.564$.
3. Mean first exit time
In this section, we derive precise estimate on the average of the first exit time $\widetilde{\tau}_{\varepsilon}$ (recall (1.19)) from the domain $\widetilde{\Omega}_{\varepsilon}$, where $\widetilde{\Omega}_{\varepsilon}$ is given by (1.18). We start by making precise in Section 3.1 the operator ~$\widetilde{\mathcal{L}}^{\varepsilon}$, then provide technical results in Section 3.2, which are used in Section 3.3 to prove Theorem 1.3.
3.1. Properties of the operator ~$\widetilde{\mathcal{L}}^{\varepsilon}$
From now on, we consider the mixed Dirichlet–Neumann Laplacian operator ~$\widetilde{\mathcal{L}}^{\varepsilon}$ on the domain $\widetilde{\Omega}_{\varepsilon}$ defined by (1.18), with domain:
(3.1) $\mathcal{D}(\widetilde{\mathcal{L}}^{\varepsilon})=\left\{u\in H^{1}(\widetilde{\Omega}_{\varepsilon}),\,\,\Delta u\in L^{2}(\widetilde{\Omega}_{\varepsilon}),\,\,\partial_{n}u_{|_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}}}=0,\,\,u_{|_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}}}=0\right\},$
where, as explained in the introduction, $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}=\cup_{k=1}^{N}\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}$ with $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}$ the connected component of $\Omega\cap(\varphi^{\varepsilon})^{-1}\{0\}$ closest to $x^{(k)}$, and $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}$ the remaining part of the boundary $\partial\widetilde{\Omega}_{\varepsilon}$; see (1.20). The fact that the definition of $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}$ makes sense is ensured by Lemma 3.2 below.
Let us emphasize that the spectral results presented in Section 2 on ${\mathcal{L}}^{\varepsilon}$ (Theorem 1.1 and Proposition 2.1) also hold true for the modified operator ~$\widetilde{\mathcal{L}}^{\varepsilon}$:
  • (R1)
    ~$\widetilde{\mathcal{L}}^{\varepsilon}$ is nonnegative self-adjoint and with compact resolvent;
  • (R2)
    There exist $c>0$ and $\varepsilon_{0}>0$ such that for any $\varepsilon\in(0,\varepsilon_{0})$, $\operatorname{dim}\operatorname{Ran}\hskip 1.9919pt\pi_{[0,c\hskip 0.28436pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon})=1$;
  • (R3)
    The second eigenvalue $\widetilde{\lambda}^{\varepsilon}_{1}$ of the operator ~$\widetilde{\mathcal{L}}^{\varepsilon}$ satisfies
    (3.2) $\lambda^{\mathbf{N}}_{1}(\widetilde{\Omega}_{\varepsilon})\leqslant\widetilde{\lambda}^{\varepsilon}_{1}\leqslant\lambda^{\mathbf{D}}_{1}(\widetilde{\Omega}_{\varepsilon}),$
    where $\lambda^{\mathbf{N}}_{1}(\widetilde{\Omega}_{\varepsilon})$ (resp. $\lambda^{\mathbf{D}}_{1}(\widetilde{\Omega}_{\varepsilon})$) is the second eigenvalue of the Neumann (resp. Dirichlet) problem of the Laplacian on the domain $\widetilde{\Omega}_{\varepsilon}$.
To prove (R2), it suffices in the proof written in Section 2.2 to replace the domain $\Omega$ with the modified domain $\widetilde{\Omega}_{\varepsilon}$, and the quasi-mode $\varphi_{\alpha}^{\varepsilon}$ in (2.10) by $\varphi_{1}^{\varepsilon}=\varphi^{\varepsilon}$, as this function satisfies by construction the required boundary conditions.
In order to extract a useful lower bound from (R3), we need a positive lower bound uniform in $\varepsilon$ on $\lambda^{\mathbf{N}}_{1}(\widetilde{\Omega}_{\varepsilon})$. This is equivalent to bounding from below the Poincaré–Wirtinger constant of the domain $\widetilde{\Omega}_{\varepsilon}$. The following lemma is a direct consequence of [44, Theorem 1.2] (see also [6]), since the family of domains $(\widetilde{\Omega}_{\varepsilon})_{\varepsilon>0}$ admits a uniform in $\varepsilon$ interior cone condition in view of (1.26).
Lemma 3.1.
Let $\lambda^{\mathbf{N}}_{1}(\widetilde{\Omega}_{\varepsilon})$ be the second eigenvalue of the Laplacian on the domain $\widetilde{\Omega}_{\varepsilon}$, with Neumann boundary conditions. Then, there exists $\underline{\lambda}^{\mathbf{N}}_{1}>0$ such that, for any $\varepsilon\in(0,\varepsilon_{0})$,
(3.3) $\underline{\lambda}^{\mathbf{N}}_{1}\leqslant\lambda^{\mathbf{N}}_{1}(\widetilde{\Omega}_{\varepsilon}).$
A consequence of (3.2) and Lemma 3.1 is that $\widetilde{\lambda}^{\varepsilon}_{1}$ admits $\underline{\lambda}^{\mathbf{N}}_{1}$ as a lower bound independently of $\varepsilon\in(0,\varepsilon_{0})$.
We conclude this section by a result ensuring that the Dirichlet region is well defined.
Lemma 3.2.
Upon further reducing $\varepsilon_{0}>0$ so that, for any $\varepsilon\in(0,\varepsilon_{0})$ and $k\in\{1,\ldots,K\}$,
(3.4) $r^{(k)}_{\varepsilon,+}\leqslant\frac{\rho_{0}}{2},\qquad K_{\varepsilon}^{(k)}-r^{(k)}_{\varepsilon,+}\left(\frac{2}{\rho_{0}}+\frac{1}{2}\right)\geqslant 0,$
the set $\Omega\cap\mathbf{B}\left(x^{(k)},r^{(k)}_{\varepsilon,+}\right)\cap\left(\varphi^{\varepsilon}\right)^{-1}[0,+\infty)$ is star shaped with respect to $x^{(k)}$ for any $k\in\{1,\ldots,K\}$, and therefore connected.
In particular, $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}$ can indeed be defined as the connected component of $\Omega\cap(\varphi^{\varepsilon})^{-1}\{0\}$ closest to $x^{(k)}$. Note that the second condition of (3.4) can indeed be satisfied since $r^{(k)}_{\varepsilon,+}=\mathrm{e}^{-C_{+}/K_{\varepsilon}^{(k)}}$ goes to zero much faster than $K_{\varepsilon}^{(k)}$.
Proof.
Fix $k\in\{1,\dots,K\}$ and $x\in\Omega\cap\mathbf{B}\left(x^{(k)},r^{(k)}_{\varepsilon,+}\right)\cap\left(\varphi^{\varepsilon}\right)^{-1}[0,+\infty)$. We prove that the segment $\{(1-t)x+tx^{(k)},\ t\in[0,1]\}$ is included in $\Omega\cap\mathbf{B}\left(x^{(k)},r^{(k)}_{\varepsilon,+}\right)\cap\left(\varphi^{\varepsilon}\right)^{-1}[0,+\infty)$. This implies that $\Omega\cap\mathbf{B}\left(x^{(k)},r^{(k)}_{\varepsilon,+}\right)\cap\left(\varphi^{\varepsilon}\right)^{-1}[0,+\infty)$ is star shaped with respect to $x^{(k)}$.
Consider the function $[0,1)\ni t\mapsto h(t)=\varphi^{\varepsilon}((1-t)x+tx^{(k)})$. It suffices to prove that $h$ is nondecreasing for $t\in[0,1)$. Indeed, if this is the case, then $h(t)\geqslant h(0)\geqslant 0$ for any $t\in[0,1)$, so that $\{(1-t)x+tx^{(k)},\ t\in[0,1]\}$ is included in $\Omega\cap\mathbf{B}\left(x^{(k)},r^{(k)}_{\varepsilon,+}\right)\cap\left(\varphi^{\varepsilon}\right)^{-1}[0,+\infty)$.
To prove that $h$ is nondecreasing, we show that its derivative is nonnegative. In view of (1.17) and (2.18), a simple computation shows that
$h^{\prime}(t)$ $=\frac{K_{\varepsilon}^{(k)}}{\sqrt{\pi}}\left(x-x^{(k)}\right)\cdot\left[\frac{x-x^{(k)}}{(1-t)|x-x^{(k)}|^{2}}-\frac{(1-t)x+tx^{(k)}}{2}\right]$
$\quad+\left(x-x^{(k)}\right)\cdot\sum_{\ell\neq k}\frac{K_{\varepsilon}^{(\ell)}}{\sqrt{\pi}}\nabla f_{\ell}\left((1-t)x+tx^{(k)}\right)$
$\geqslant\frac{K_{\varepsilon}^{(k)}}{\sqrt{\pi}}\left(\frac{1}{1-t}-\frac{1}{2}\right)-r^{(k)}_{\varepsilon,+}\sum_{\ell\neq k}\frac{K_{\varepsilon}^{(\ell)}}{\sqrt{\pi}}\left|\nabla f_{\ell}\left((1-t)x+tx^{(k)}\right)\right|,$
where we used $|(1-t)x+tx^{(k)}|\leqslant 1$ for all $t\in[0,1]$ and $|x-x^{(k)}|\leqslant r^{(k)}_{\varepsilon,+}\leqslant 1$. For $\ell\neq k$, it holds, for any $t\in[0,1]$,
$\left|\nabla f_{\ell}\left((1-t)x+tx^{(k)}\right)\right|$ $\leqslant\frac{1}{|(1-t)(x-x^{(k)})+(x^{(k)}-x^{(\ell)})|}+\frac{1}{2}\leqslant\frac{1}{\rho_{0}-r^{(k)}_{\varepsilon,+}}+\frac{1}{2},$
since $|(1-t)(x-x^{(k)})+(x^{(k)}-x^{(\ell)})|\geqslant|x^{(k)}-x^{(\ell)}|-(1-t)|x-x^{(k)}|\geqslant\rho_{0}-(1-t)r^{(k)}_{\varepsilon,+}\geqslant\rho_{0}-r^{(k)}_{\varepsilon,+}>0$ in view of (3.4). Thus, using (1.11),
$h^{\prime}(t)$ $\geqslant\frac{K_{\varepsilon}^{(k)}}{2\sqrt{\pi}}-r^{(k)}_{\varepsilon,+}\frac{\overline{K}_{\varepsilon}}{\sqrt{\pi}}\left(\frac{1}{\rho_{0}-r^{(k)}_{\varepsilon,+}}+\frac{1}{2}\right)\geqslant\frac{1}{2\sqrt{\pi}}\left[K_{\varepsilon}^{(k)}-r^{(k)}_{\varepsilon,+}\left(\frac{1}{\rho_{0}-r^{(k)}_{\varepsilon,+}}+\frac{1}{2}\right)\right]\geqslant 0,$
where the last inequality follows from the two conditions in (3.4). This allows to conclude the proof. ∎
3.2. Useful technical results
Let us now state several lemmas that will be useful to prove Theorem 1.3. The following lemma concerns the function ${\varphi}^{\varepsilon}$.
Lemma 3.3.
Let ${\varphi}^{\varepsilon}$ be the function defined in (1.17). Then, ${\varphi}^{\varepsilon}$ belongs to $\mathcal{D}(\widetilde{\mathcal{L}}^{\varepsilon})$ and satisfies
(3.5) $\Delta{\varphi}^{\varepsilon}=\frac{\overline{K}_{\varepsilon}}{\sqrt{\pi}}\text{ in }\widetilde{\Omega}_{\varepsilon},$
where $\overline{K}_{\varepsilon}$ is defined by (1.10).
Proof.
The function ${\varphi}^{\varepsilon}$ belongs to $C^{\infty}(\widetilde{\Omega}_{\varepsilon})$ and satisfies ${\varphi}^{\varepsilon}=0$ on $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}$ by construction. The equality (2.14) implies moreover that $\nabla{\varphi}^{\varepsilon}\cdot\vec{n}=0$ on $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}$, so that ${\varphi}^{\varepsilon}\in\mathcal{D}(\widetilde{\mathcal{L}}^{\varepsilon})$. Finally, (3.5) is a consequence of (1.17) and (2.15), which concludes the proof. ∎
We can then write the following estimate on the quasi-mode ${\varphi}^{\varepsilon}$ for ~$\widetilde{\mathcal{L}}^{\varepsilon}$, giving in particular bounds on the distance to normalized elements in the first eigenspace ${\rm Ran}(\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}))$.
Proposition 3.4.
There exists $C\in\mathbb{R}_{+}$ such that, for any $\varepsilon\in(0,\varepsilon_{0})$,
(3.6) $\left\|{\varphi}^{\varepsilon}-\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\varphi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}\leqslant C\overline{K}_{\varepsilon}.$
In particular, upon possibly reducing $\varepsilon_{0}$, it holds $\left\|\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\varphi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}\geqslant 1/2$, so that the first eigenfunction of ~$\widetilde{\mathcal{L}}^{\varepsilon}$ can be obtained from $\varphi^{\varepsilon}$ as
(3.7) $\widetilde{u}^{\varepsilon}_{0}=\frac{\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\varphi}^{\varepsilon}}{\left\|\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\varphi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}}.$
Moreover,
(3.8) $\left\|{\varphi}^{\varepsilon}-\widetilde{u}^{\varepsilon}_{0}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}\leqslant C\overline{K}_{\varepsilon}.$
This shows that ${\varphi}^{\varepsilon}$ is a quasi-mode (1.17) associated with the eigenvalue $\widetilde{\lambda}^{\varepsilon}_{0}$.
Note that the equality (3.7) fixes the sign convention for $\widetilde{u}^{\varepsilon}_{0}$.
Proof.
We start by proving (3.6). Note first that $\sigma(\widetilde{\mathcal{L}}^{\varepsilon})\cap B(0,\underline{\lambda}^{\mathbf{N}}_{1}/2)=\{\widetilde{\lambda}_{0}^{\varepsilon}\}$ thanks to (R2) and the lower bound on the second eigenvalue $\tilde{\lambda}_{0}^{\varepsilon}$ of ~$\widetilde{\mathcal{L}}^{\varepsilon}$ obtained in Lemma 3.1. Since ${\varphi}^{\varepsilon}\in\mathcal{D}(\widetilde{\mathcal{L}}^{\varepsilon})$ (see Lemma 3.3), we obtain that
$\left(1-\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon})\right){\varphi}^{\varepsilon}$ $=\frac{1}{2\pi\mathrm{i}}\int_{\mathscr{C}\left(0,\,{\underline{\lambda}^{\mathbf{N}}_{1}}/{2}\right)}\left(z^{-1}-(z-\widetilde{\mathcal{L}}^{\varepsilon})^{-1}\right){\varphi}^{\varepsilon}\,\mathrm{d}z$
$=-\frac{1}{2\pi\mathrm{i}}\int_{\mathscr{C}\left(0,\,{\underline{\lambda}^{\mathbf{N}}_{1}}/{2}\right)}z^{-1}(z-\widetilde{\mathcal{L}}^{\varepsilon})^{-1}\widetilde{\mathcal{L}}^{\varepsilon}{\varphi}^{\varepsilon}\,\mathrm{d}z,$
where $\mathscr{C}\left(0,\,{\underline{\lambda}^{\mathbf{N}}_{1}}/{2}\right)\subset\mathbb{C}$ is the circle of radius ${\underline{\lambda}^{\mathbf{N}}_{1}}/{2}$ centered at 00. Using again the upper bound on $\tilde{\lambda}_{0}^{\varepsilon}$ and the lower bound on $\tilde{\lambda}_{1}^{\varepsilon}$, and classical resolvent estimates, we obtain that
$\forall z\in\mathscr{C}\left(0,\,{\underline{\lambda}^{\mathbf{N}}_{1}}/{2}\right),\qquad\left\|(z-\widetilde{\mathcal{L}}^{\varepsilon})^{-1}\right\|_{\mathcal{B}({L^{2}(\widetilde{\Omega}_{\varepsilon})})}\leqslant\frac{C}{{\underline{\lambda}^{\mathbf{N}}_{1}}},$
where $C>0$ is independent of $\varepsilon$. Here and in the following, $\mathcal{B}(L^{2}(\widetilde{\Omega}_{\varepsilon}))$ is the Banach space of bounded operators from $L^{2}(\widetilde{\Omega}_{\varepsilon})$ to $L^{2}(\widetilde{\Omega}_{\varepsilon})$, with $\left\|\cdot\right\|_{\mathcal{B}({L^{2}(\widetilde{\Omega}_{\varepsilon})})}$ the associated operator norm. As a consequence, using Lemma 3.3, for $\varepsilon$ small enough, we have
$\left\|\left(1-\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon})\right){\varphi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}\leqslant\frac{C}{(\underline{\lambda}^{\mathbf{N}}_{1})^{2}}\left\|\widetilde{\mathcal{L}}^{\varepsilon}{\varphi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}\leqslant C\overline{K}_{\varepsilon},$
where $C>0$ is independent of $\varepsilon$.
An immediate consequence of (3.6) is that
(3.9) $\left\|\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\varphi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}^{2}=\left\|{\varphi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}^{2}-\left\|{\varphi}^{\varepsilon}-\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\varphi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}^{2}=1+\mathrm{O}(\overline{K}_{\varepsilon}),$
where we used the estimate (2.26) on $\|{\varphi}^{\varepsilon}\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}$ (with $\alpha=1$).
Let us finally turn to (3.8). Using the triangle inequality, then (3.6) and (3.9),
$\left\|{\varphi}^{\varepsilon}-\widetilde{u}^{\varepsilon}_{0}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}$ $\leqslant\left\|{\varphi}^{\varepsilon}-\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\varphi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}+\left\|\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\varphi}^{\varepsilon}-\frac{\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\varphi}^{\varepsilon}}{\|\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\varphi}^{\varepsilon}\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}$
$=\left\|{\varphi}^{\varepsilon}-\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\varphi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}+1-\left\|\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\varphi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}=\mathrm{O}(\overline{K}_{\varepsilon}),$
which allows to conclude the proof of Proposition 3.4. ∎
3.3. Proof of Theorem 1.3
We can finally write the proof of Theorem 1.3. It is convenient to introduce the normalized function
(3.10) ${\psi}^{\varepsilon}=\frac{{\varphi}^{\varepsilon}}{\|{\varphi}^{\varepsilon}\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}}.$
We claim that
(3.11) $\left\|\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\psi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}=1+\mathrm{O}\left(\overline{K}^{2}_{\varepsilon}\right),$
and
(3.12) $\left\|\psi^{\varepsilon}-\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\psi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}=\mathrm{O}\left(\overline{K}_{\varepsilon}\right),$
Indeed, using (2.26), (3.6) and the fact that $\|{\psi}^{\varepsilon}\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}=1$,
$\left\|\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\psi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}^{2}$ $=\|{\psi}^{\varepsilon}\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}^{2}-\left\|\left(1-\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon})\right){\psi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}^{2}$
$=1-\frac{\left\|\left(1-\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon})\right){\varphi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}^{2}}{\|\varphi^{\varepsilon}\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}^{2}}=1+\frac{\mathrm{O}\left(\overline{K}_{\varepsilon}^{2}\right)}{1+\mathrm{O}(\overline{K}_{\varepsilon})},$
which leads to (3.11). The estimate (3.12) then follows from the equality
$\left\|\psi^{\varepsilon}-\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\psi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}^{2}=1-\left\|\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\psi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}^{2}.$
We can now turn to the estimation of the first eigenvalue. Noting that (3.7) holds with $\varphi^{\varepsilon}$ replaced by $\psi^{\varepsilon}$, and using next (3.11),
(3.13) $\begin{split}\widetilde{\lambda}_{0}^{\varepsilon}=-\left\langle\Delta\widetilde{u}^{\varepsilon}_{0},\widetilde{u}^{\varepsilon}_{0}\right\rangle_{L^{2}(\widetilde{\Omega}_{\varepsilon})}=-\left\langle\Delta{\psi}^{\varepsilon},\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\psi}^{\varepsilon}\right\rangle_{L^{2}(\widetilde{\Omega}_{\varepsilon})}\left(1+\mathrm{O}\left(\overline{K}^{2}_{\varepsilon}\right)\right),\end{split}$
since the orthogonal projector $\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon})$ and ~$\widetilde{\mathcal{L}}^{\varepsilon}$ commute on $\mathcal{D}(\widetilde{\mathcal{L}}^{\varepsilon})$ and ${\psi}^{\varepsilon}\in\mathcal{D}(\widetilde{\mathcal{L}}^{\varepsilon})$. In addition,
$\left\langle\Delta{\psi}^{\varepsilon},\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\psi}^{\varepsilon}\right\rangle_{L^{2}(\widetilde{\Omega}_{\varepsilon})}=\left\langle\Delta{\psi}^{\varepsilon},{\psi}^{\varepsilon}\right\rangle_{L^{2}(\widetilde{\Omega}_{\varepsilon})}-\left\langle\Delta{\psi}^{\varepsilon},\left(1-\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon})\right){\psi}^{\varepsilon}\right\rangle_{L^{2}(\widetilde{\Omega}_{\varepsilon})}.$
For the second term on the right hand side of the previous equality,
[…]
where we used (3.12) for the first factor, and Lemma 3.3 together with the estimate (2.26) (with $\alpha=1$) for the second one. For the first term, one has
$\left\langle\Delta{\psi}^{\varepsilon},{\psi}^{\varepsilon}\right\rangle_{L^{2}(\widetilde{\Omega}_{\varepsilon})}=\left\langle\Delta{\varphi}^{\varepsilon},{\varphi}^{\varepsilon}\right\rangle_{L^{2}(\widetilde{\Omega}_{\varepsilon})}\left(1+\mathrm{O}\left(\overline{K}_{\varepsilon}\right)\right),$
where we used again the estimate (2.26) on $\|{\varphi}^{\varepsilon}\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}$. Moreover, using (1.17) and (3.5),
$\left\langle\Delta{\varphi}^{\varepsilon},{\varphi}^{\varepsilon}\right\rangle_{L^{2}(\widetilde{\Omega}_{\varepsilon})}$
$\qquad=-\frac{\overline{K}_{\varepsilon}}{\pi}\int_{\widetilde{\Omega}_{\varepsilon}}\mathrm{d}x-\frac{\overline{K}_{\varepsilon}}{\pi}\sum_{k=1}^{N}{K}^{(k)}_{\varepsilon}\int_{\widetilde{\Omega}_{\varepsilon}}\text{log}\hskip 2.84544pt|x-x^{(k)}|\,\mathrm{d}x-\frac{\overline{K}^{2}_{\varepsilon}}{\pi}\int_{\widetilde{\Omega}_{\varepsilon}}\left(\frac{1-|x|^{2}}{4}\right)\,\mathrm{d}x$
$\qquad=-\overline{K}_{\varepsilon}+\frac{\overline{K}_{\varepsilon}}{\pi}\left|\Omega\backslash\widetilde{\Omega}_{\varepsilon}\right|-\frac{\overline{K}_{\varepsilon}}{\pi}\sum_{k=1}^{N}{K}^{(k)}_{\varepsilon}\int_{\widetilde{\Omega}_{\varepsilon}}\text{log}\hskip 2.84544pt|x-x^{(k)}|\,\mathrm{d}x-\frac{\overline{K}^{2}_{\varepsilon}}{\pi}\int_{\widetilde{\Omega}_{\varepsilon}}\left(\frac{1-|x|^{2}}{4}\right)\,\mathrm{d}x.$
In view of Lemma 1.2, it holds $|\Omega\backslash\widetilde{\Omega}_{\varepsilon}|\leqslant\pi\displaystyle\sum_{k=1}^{N}\left({r^{(k)}_{\varepsilon,+}}\right)^{2}=\mathrm{O}(\overline{K}_{\varepsilon})$. Concerning the second term,
$\frac{\overline{K}_{\varepsilon}}{\pi}\sum_{k=1}^{N}{K}^{(k)}_{\varepsilon}\int_{\widetilde{\Omega}_{\varepsilon}}\text{log}\hskip 2.84544pt|x-x^{(k)}|\,\mathrm{d}x=\mathrm{O}\left(\overline{K}^{2}_{\varepsilon}\right).$
Finally,
$\frac{\overline{K}^{2}_{\varepsilon}}{\pi}\left|\int_{\widetilde{\Omega}_{\varepsilon}}\left(\frac{1-|x|^{2}}{4}\right)\,\mathrm{d}x\right|\leqslant\frac{\overline{K}^{2}_{\varepsilon}}{4}.$
The proof of Theorem 1.3 follows by gathering the above estimates in (3.13).
4. Law of the first exit point
We study in this section the law of the first exit point $X_{\widetilde{\tau_{\varepsilon}}}$ from the domain $\widetilde{\Omega}_{\varepsilon}$, where $\widetilde{\Omega}_{\varepsilon}$ and $\widetilde{\tau_{\varepsilon}}$ are respectively defined in (1.18) and (1.19). Recall that the law of $X_{\widetilde{\tau_{\varepsilon}}}$ is given by (see (1.8))
(4.1) $\frac{\partial_{n}\widetilde{u}_{0}^{\varepsilon}(x)\,\sigma(dx)}{\displaystyle\int_{\partial\widetilde{\Omega}_{\varepsilon}}\partial_{n}\widetilde{u}_{0}^{\varepsilon}(y)\,\sigma(\mathrm{d}y)},$
where $\sigma$ here denotes (with a slight abuse of notation) the Lebesgue surface measure on $\partial\widetilde{\Omega}_{\varepsilon}$.
We now prove several lemmas that lead to the result of Theorem 1.4. We begin with two results on the $L^{2}(\widetilde{\Omega}_{\varepsilon})$-normalized principal eigenfunction $\widetilde{u}^{\varepsilon}_{0}$ of ~$\widetilde{\mathcal{L}}^{\varepsilon}$. Let us recall that the sign of this eigenfunction is fixed in order for (3.7) to hold.
Lemma 4.1.
The principal eigenfunction $\widetilde{u}^{\varepsilon}_{0}$ satisfies, in the limit $\varepsilon\rightarrow 0$,
(4.2) $\int_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}}\nabla\widetilde{u}^{\varepsilon}_{0}\cdot\vec{n}\,\mathrm{d}\sigma=\sqrt{\pi}\ \overline{K}_{\varepsilon}\left(1+\mathrm{O}\left(\overline{K}_{\varepsilon}\right)\right).$
Proof.
Recall that $\widetilde{\lambda}_{0}^{\varepsilon}=\|\nabla\widetilde{u}^{\varepsilon}_{0}\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}^{2}=\left\langle\nabla\widetilde{u}^{\varepsilon}_{0},\nabla\widetilde{u}^{\varepsilon}_{0}\right\rangle_{L^{2}(\widetilde{\Omega}_{\varepsilon})}$. Using (3.7), the fact that
$\nabla\pi_{[0,c\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon})=\pi_{[0,c\overline{K}_{\varepsilon}]}(\Delta^{(1)})\nabla,$
where $\Delta^{(1)}$ here denotes the Laplacian on $1$-forms with mixed tangential-normal boundary conditions (see (A.10) and Proposition A.3, applied with $\Omega=\widetilde{\Omega}_{\varepsilon}$, $\Gamma^{\varepsilon}_{\mathbf{D}}=\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}$ and $\Gamma^{\varepsilon}_{\mathbf{N}}=\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}$), and $\pi_{[0,c\overline{K}_{\varepsilon}]}(\Delta^{(1)})\nabla\widetilde{u}^{\varepsilon}_{0}=\nabla\widetilde{u}^{\varepsilon}_{0}$, we get
$\widetilde{\lambda}_{0}^{\varepsilon}$ $=\frac{\left\langle\nabla{\varphi}^{\varepsilon},\nabla\widetilde{u}^{\varepsilon}_{0}\right\rangle_{L^{2}(\widetilde{\Omega}_{\varepsilon})}}{\left\|\pi_{[0,c\hskip 0.56917pt\overline{K}_{\varepsilon}]}(\widetilde{\mathcal{L}}^{\varepsilon}){\varphi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}}.$
Using (3.9) and Theorem 1.3, it follows that
(4.3) $\left\langle\nabla{\varphi}^{\varepsilon},\nabla\widetilde{u}^{\varepsilon}_{0}\right\rangle_{L^{2}(\widetilde{\Omega}_{\varepsilon})}=\overline{K}_{\varepsilon}\left(1+\mathrm{O}\left(\overline{K}_{\varepsilon}\right)\right).$
Besides, using the Green formula,
$\left\langle\nabla{\varphi}^{\varepsilon},\nabla\widetilde{u}^{\varepsilon}_{0}\right\rangle_{L^{2}(\widetilde{\Omega}_{\varepsilon})}$ $=\left\langle\nabla\left(\frac{1}{\sqrt{\pi}}+{\varphi}^{\varepsilon}\right),\nabla\widetilde{u}^{\varepsilon}_{0}\right\rangle_{L^{2}(\widetilde{\Omega}_{\varepsilon})}$
$=-\left\langle\left(\frac{1}{\sqrt{\pi}}+{\varphi}^{\varepsilon}\right),\Delta\widetilde{u}^{\varepsilon}_{0}\right\rangle_{L^{2}(\widetilde{\Omega}_{\varepsilon})}+\int_{\partial\widetilde{\Omega}_{\varepsilon}}\left(\frac{1}{\sqrt{\pi}}+{\varphi}^{\varepsilon}\right)\nabla\widetilde{u}^{\varepsilon}_{0}\cdot\vec{n}.$
Since ${\varphi}^{\varepsilon}=0$ on $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}$ and $\nabla{\widetilde{u}}^{\varepsilon}_{0}\cdot\vec{n}=0$ on $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}$, the last term in the previous equality is equal to
(4.4) $\int_{\partial\widetilde{\Omega}_{\varepsilon}}\left(\frac{1}{\sqrt{\pi}}+{\varphi}^{\varepsilon}\right)\nabla\widetilde{u}^{\varepsilon}_{0}\cdot\vec{n}=\frac{1}{\sqrt{\pi}}\int_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}}\nabla\widetilde{u}^{\varepsilon}_{0}\cdot\vec{n}.$
Moreover,
$\left|\left\langle\left(\frac{1}{\sqrt{\pi}}+{\varphi}^{\varepsilon}\right),\Delta\widetilde{u}^{\varepsilon}_{0}\right\rangle_{L^{2}(\widetilde{\Omega}_{\varepsilon})}\right|$ $\leqslant\left\|\frac{1}{\sqrt{\pi}}+{\varphi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}\|\Delta\widetilde{u}^{\varepsilon}_{0}\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}$
$=\widetilde{\lambda}_{0}^{\varepsilon}\left\|\frac{1}{\sqrt{\pi}}+{\varphi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}\leqslant C\overline{K}^{2}_{\varepsilon},$
where we used Theorem 1.1 and
(4.5) $\left\|\frac{1}{\sqrt{\pi}}+{\varphi}^{\varepsilon}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}=\mathrm{O}(\overline{K}_{\varepsilon})$
which follows from (2.29) (with $\alpha=1$). This concludes the proof of (4.2). ∎
Lemma 4.2.
There exists $C\in\mathbb{R}_{+}$ such that, for any $k\in\{1,\dots,N\}$ and for any $\varepsilon\in(0,\varepsilon_{0})$,
(4.6) $\left|\int_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}}\nabla(\widetilde{u}_{0}^{\varepsilon}-{\varphi}^{\varepsilon})\cdot\vec{n}\,\mathrm{d}\sigma\right|\leqslant C\overline{K}^{3/2}_{\varepsilon},$
where $\widetilde{u}_{0}^{\varepsilon}$ is the $L^{2}(\widetilde{\Omega}_{\varepsilon})$-normalized principal eigenfunction of ~$\widetilde{\mathcal{L}}^{\varepsilon}$ (with the appropriate sign convention such that (3.7) holds) and ${\varphi}^{\varepsilon}$ is the quasi-mode defined in (1.17).
Proof.
Let us introduce ${v}^{\varepsilon}:=\widetilde{u}_{0}^{\varepsilon}-{\varphi}^{\varepsilon}$. We start by writing preliminary estimates on $\Delta v^{\varepsilon}$ and $\nabla v^{\varepsilon}$, and then relate with a Green formula the integral on $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}$ in (4.6) to integrals over $\widetilde{\Omega}_{\varepsilon}$ involving $\Delta v^{\varepsilon}$ and $\nabla v^{\varepsilon}$.
We first claim that
(4.7) $\|\Delta{v}^{\varepsilon}\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}=\mathrm{O}\left(\overline{K}^{2}_{\varepsilon}\right).$
Indeed, in view of (1.24) and Lemma 3.3,
$\Delta{v}^{\varepsilon}$ $=-\widetilde{\lambda}_{0}^{\varepsilon}\widetilde{u}_{0}^{\varepsilon}-\frac{\overline{K}_{\varepsilon}}{\sqrt{\pi}}=\left(-\widetilde{\lambda}_{0}^{\varepsilon}+\overline{K}_{\varepsilon}\right)\widetilde{u}_{0}^{\varepsilon}+\overline{K}_{\varepsilon}\left({\varphi}^{\varepsilon}-\widetilde{u}_{0}^{\varepsilon}\right)-\overline{K}_{\varepsilon}\left(\frac{1}{\sqrt{\pi}}+{\varphi}^{\varepsilon}\right).$
Therefore, Theorem 1.3, (3.8), and (4.5) imply that (4.7) holds.
We next prove that
(4.8) $\|\nabla{v}^{\varepsilon}\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}=\mathrm{O}\left(\overline{K}^{3/2}_{\varepsilon}\right).$
Since ${v}^{\varepsilon}=0$ on $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}$ and $\nabla{v}^{\varepsilon}\cdot\vec{n}=0$ on $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}$, by integration by parts,
$\int_{\widetilde{\Omega}_{\varepsilon}}|\nabla{v}^{\varepsilon}|^{2}=\int_{\widetilde{\Omega}_{\varepsilon}}\left(\Delta{v}^{\varepsilon}\right){v}^{\varepsilon}.$
Now, using a Cauchy–Schwarz inequality, (3.8), and (4.7), we obtain that, for $\varepsilon$ small enough, enough,
$\left|\int_{\widetilde{\Omega}_{\varepsilon}}\left(\Delta{v}^{\varepsilon}\right){v}^{\varepsilon}\right|\leqslant\|\Delta{v}^{\varepsilon}\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}\|{v}^{\varepsilon}\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}\leqslant C\overline{K}^{3}_{\varepsilon},$
which yields (4.8).
We now fix $k\in\{1,\dots,N\}$ and relate the integral on $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}$ in (4.6) to integrals over $\widetilde{\Omega}_{\varepsilon}$ involving $\Delta v^{\varepsilon}$ and $\nabla v^{\varepsilon}$. We introduce to this end a smooth cut-off function $\chi^{\ell}_{k}:{\Omega}\rightarrow[0,1]$, defined using some parameter $10<\ell<$ 1, which satisfies
$\chi^{\ell}_{k}(x)=1\textrm{ on }{\mathbf{B}}\left(x^{(k)},\frac{\ell}{2}\right)\cap\Omega,\qquad\chi^{\ell}_{k}(x)=0\textrm{ on }\Omega\backslash{\mathbf{B}}\left(x^{(k)},{\ell}\right).$
The parameter $\ell\in(0,\rho_{0})$ is chosen such that, for any $\varepsilon\in(0,\varepsilon_{0})$,
$\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}\subset{\mathbf{B}}\left(x^{(k)},\frac{\ell}{2}\right)\cap\Omega,\qquad\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k^{\prime}}}\cap\mathbf{B}\left(x^{(k)},\ell\right)=\emptyset\textrm{ for }k^{\prime}\neq k.$
Since ${v}^{\varepsilon}\in\mathcal{D}(\widetilde{\mathcal{L}}^{\varepsilon})$ (because $\widetilde{u}_{0}^{\varepsilon}$ and ${\varphi}^{\varepsilon}$ belong to $\mathcal{D}(\widetilde{\mathcal{L}}^{\varepsilon})$, see Lemma 3.3), $\nabla{v}^{\varepsilon}$ belongs to the vector space $\mathcal{D}(Q^{(1)})$ defined in (A.11) with $\Omega$, $\Gamma^{\varepsilon}_{\mathbf{D}}$, $\Gamma^{\varepsilon}_{\mathbf{N}}$ replaced by $\widetilde{\Omega}_{\varepsilon}$, $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}$, $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}$. Moreover, recall that $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}$ and $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}$ meet at an angle equal to $\frac{\pi}{2}$ for all $k\in\{1,\dots,N\}$ (see (1.26); as well as [33, Definition 31] for more details about the angle between two hypersurfaces). Therefore, [33, Proposition 32], which is a direct consequence of [25, 17], implies that
$\partial_{n}{v}^{\varepsilon}\in L^{2}(\partial\widetilde{\Omega}_{\varepsilon}).$
Then, using an integration by parts, we have
(4.9) $\int_{\widetilde{\Omega}_{\varepsilon}}\Delta{v}^{\varepsilon}\chi^{\ell}_{k}=-\int_{\widetilde{\Omega}_{\varepsilon}}\nabla{v}^{\varepsilon}\cdot\nabla\chi^{\ell}_{k}+\int_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}}\partial_{n}{v}^{\varepsilon}\chi^{\ell}_{k}\,\mathrm{d}\sigma,$
where we used that $\partial_{n}{v}^{\varepsilon}=0$ on $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}$ and $\chi^{\ell}_{k}=0$ on $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k^{\prime}}}$ for $k^{\prime}\neq k$. Using a Cauchy–Schwarz inequality and (4.7) we obtain that, for $\varepsilon$ small enough,
(4.10) $\left|\int_{\widetilde{\Omega}_{\varepsilon}}\Delta{v}^{\varepsilon}\chi^{\ell}_{k}\right|\leqslant\|\Delta{v}^{\varepsilon}\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}\left\|\chi^{\ell}_{k}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}\leqslant C\overline{K}^{2}_{\varepsilon}.$
Let us next deal with the first term on the right hand side of (4.9). A Cauchy–Schwarz inequality and (4.8) imply that, for $\varepsilon$ small enough,
(4.11) $\left|\int_{\widetilde{\Omega}_{\varepsilon}}\nabla{v}^{\varepsilon}\cdot\nabla\chi^{\ell}_{k}\,\mathrm{d}x\right|\leqslant\|\nabla{v}^{\varepsilon}\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}\left\|\nabla\chi^{\ell}_{k}\right\|_{L^{2}(\widetilde{\Omega}_{\varepsilon})}\leqslant C\overline{K}^{3/2}_{\varepsilon}.$
Finally, using (4.9), (4.10) and (4.11), we obtain that, for $\varepsilon$ small enough,
$\int_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}}\partial_{n}{v}^{\varepsilon}\chi^{\ell}_{k}\,\mathrm{d}\sigma=\int_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}}\partial_{n}{v}^{\varepsilon}\,\mathrm{d}\sigma=\mathrm{O}\left(\overline{K}^{3/2}_{\varepsilon}\right),$
where we used that $\chi^{\ell}_{k}=1$ on $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}$. This concludes the proof of (4.6). ∎
The next two lemmas provide explicit computations on the exit flux for the quasimode $\varphi^{\varepsilon}$.
Lemma 4.3.
Let $k\in\{1,\dots,N\}$. Then, when $\varepsilon\rightarrow 0$,
(4.12) $\int_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}}\nabla{\varphi}^{\varepsilon}\cdot\vec{n}\,\mathrm{d}\sigma=-\int_{\Omega\cap\mathscr{C}\left(x^{(k)},r^{(k)}_{\varepsilon,+}\right)}\nabla{\varphi}^{\varepsilon}\cdot\vec{n}\,\mathrm{d}\sigma+\mathrm{O}\left(\overline{K}^{2}_{\varepsilon}\right),$
where ${\varphi}^{\varepsilon}$ is the quasi-mode defined in (1.17), the normal vector on the left-hand side of (4.12) is the unit normal vector on $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}$ which points outward $\widetilde{\Omega}_{\varepsilon}$, and the normal vector on the right-hand side corresponds to the outward unit normal to $\mathbf{B}\left(x^{(k)},r^{(k)}_{\varepsilon,+}\right)$.
Proof.
Let $k\in\{1,\dots,N\}$. Let us introduce (see Figure 6)
(4.13) $\mathcal{S}^{\varepsilon}:=\widetilde{\Omega}_{\varepsilon}\cap{\mathbf{B}}\left(x^{(k)},r^{(k)}_{\varepsilon,+}\right).$
[…]
Figure 6. The domain $\mathcal{S}^{\varepsilon}$ in the neighborhood of $x^{(k)}$.
Using Stokes’ formula on $\mathcal{S}^{\varepsilon}$ (which is licit since ${\varphi}^{\varepsilon}\in C^{\infty}(\widetilde{\Omega}_{\varepsilon})$),
(4.14) $\begin{split}\int_{\mathcal{S}^{\varepsilon}}\Delta{\varphi}^{\varepsilon}&=\int_{\partial\mathcal{S}^{\varepsilon}}\nabla{\varphi}^{\varepsilon}\cdot\vec{n}\,\mathrm{d}\sigma=\int_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}}\nabla{\varphi}^{\varepsilon}\cdot\vec{n}\,\mathrm{d}\sigma+\int_{\Omega\cap\mathscr{C}\left(x^{(k)},r^{(k)}_{\varepsilon,+}\right)}\nabla{\varphi}^{\varepsilon}\cdot\vec{n}\,\mathrm{d}\sigma,\end{split}$
where we used that $\nabla{\varphi}^{\varepsilon}\cdot\vec{n}=0$ on $\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}$. Moreover, using Lemma 3.3,
(4.15) $\left|\int_{\mathcal{S}^{\varepsilon}}\Delta{\varphi}^{\varepsilon}\right|=\frac{\overline{K}_{\varepsilon}}{\sqrt{\pi}}\left|\mathcal{S}^{\varepsilon}\right|\leqslant\frac{\overline{K}_{\varepsilon}}{\sqrt{\pi}}\left|{\mathbf{B}}\left(x^{(k)},r^{(k)}_{\varepsilon,+}\right)\right|=\overline{K}_{\varepsilon}\sqrt{\pi}\left(r^{(k)}_{\varepsilon,+}\right)^{2}=\mathrm{O}\left(\overline{K}^{2}_{\varepsilon}\right),$
which allows to conclude the proof of (4.12). ∎
The first term on the right hand side of (4.12) is estimated in the following lemma.
Lemma 4.4.
Let $k\in\{1,\dots,N\}$. Then, when $\varepsilon\rightarrow 0$,
(4.16) $\int_{\Omega\cap\mathscr{C}\left(x^{(k)},r^{(k)}_{\varepsilon,+}\right)}\nabla{\varphi}^{\varepsilon}\cdot\vec{n}\,\mathrm{d}\sigma=-\sqrt{\pi}{K}^{(k)}_{\varepsilon}+\mathrm{O}\left(\overline{K}^{2}_{\varepsilon}\right),$
where ${\varphi}^{\varepsilon}$ is the quasi-mode defined in (1.17).
Proof.
Let us introduce the short-hand notation $C^{(k)}_{\varepsilon}=\Omega\cap\mathscr{C}\left(x^{(k)},r^{(k)}_{\varepsilon,+}\right)$. From the definition (1.17), we have
(4.17) $\begin{split}\int_{C^{(k)}_{\varepsilon}}\nabla{\varphi}^{\varepsilon}\cdot\vec{n}\,\mathrm{d}\sigma=-\frac{K^{(k)}_{\varepsilon}}{\sqrt{\pi}}\int_{C^{(k)}_{\varepsilon}}\nabla f_{k}\cdot\vec{n}\,\mathrm{d}\sigma-\frac{1}{\sqrt{\pi}}\sum_{\begin{subarray}{c}k^{\prime}=1\\ k^{\prime}\neq k\end{subarray}}^{N}K^{(k^{\prime})}_{\varepsilon}\int_{C^{(k)}_{\varepsilon}}\nabla f_{k^{\prime}}\cdot\vec{n}\,\mathrm{d}\sigma.\end{split}$
In view of (2.18), the first integral on the right hand side of (4.17) reads
(4.18) $\begin{split}-\frac{K^{(k)}_{\varepsilon}}{\sqrt{\pi}}\int_{C^{(k)}_{\varepsilon}}\nabla f_{k}\cdot\vec{n}\,\mathrm{d}\sigma=-\frac{K^{(k)}_{\varepsilon}}{\sqrt{\pi}}\int_{C^{(k)}_{\varepsilon}}\frac{x-x^{(k)}}{|x-x^{(k)}|^{2}}\cdot\vec{n}\,\mathrm{d}\sigma+\frac{K^{(k)}_{\varepsilon}}{\sqrt{\pi}}\int_{C^{(k)}_{\varepsilon}}\frac{x}{2}\cdot\vec{n}\,\mathrm{d}\sigma.\end{split}$
As $\vec{n}(x)=\frac{x-x^{(k)}}{|x-x^{(k)}|}$ on $\mathscr{C}\left(x^{(k)},r^{(k)}_{\varepsilon,+}\right)$ (see Figure 6) and thus
$\int_{C^{(k)}_{\varepsilon}}\frac{x-x^{(k)}}{|x-x^{(k)}|^{2}}\cdot\vec{n}\,\mathrm{d}\sigma=\int_{C^{(k)}_{\varepsilon}}\frac{1}{|x-x^{(k)}|}\,\mathrm{d}\sigma=\pi-\theta_{\varepsilon},$
where $\theta_{\varepsilon}=\mathrm{arcsin}(r^{(k)}_{\varepsilon,+}/2)$. Concerning the second integral on the right hand side of (4.18), using that $|x|\leqslant 1$ for $x\in C^{(k)}_{\varepsilon}$ we obtain
$\left|\int_{C^{(k)}_{\varepsilon}}\frac{x}{2}\cdot\vec{n}\,\mathrm{d}\sigma\right|\leqslant\frac{1}{2}\left|C^{(k)}_{\varepsilon}\right|=\frac{1}{2}r^{(k)}_{\varepsilon,+}(\pi-\theta_{\varepsilon}).$
Then, for $\varepsilon$ small enough,
(4.19) $-\frac{K^{(k)}_{\varepsilon}}{\sqrt{\pi}}\int_{C^{(k)}_{\varepsilon}}\nabla f_{k}\cdot\vec{n}\,\mathrm{d}\sigma=-{K^{(k)}_{\varepsilon}}{\sqrt{\pi}}+\mathrm{O}\left(\overline{K}^{2}_{\varepsilon}\right).$
Let us now deal with the second integral on the right hand side of (4.17). Note that, in view of (1.2), there exists $\rho>0$ independent of $\varepsilon$ such that, for $\varepsilon$ small enough,
(4.20) $\forall x\in C^{(k)}_{\varepsilon},\quad\forall k^{\prime}\neq k,\qquad|x-x^{(k^{\prime})}|\geqslant\rho.$
Using again (2.18),
$\int_{C^{(k)}_{\varepsilon}}\nabla f_{k^{\prime}}\cdot\vec{n}\,\mathrm{d}\sigma=\int_{C^{(k)}_{\varepsilon}}\frac{x-x^{(k^{\prime})}}{|x-x^{(k^{\prime})}|^{2}}\cdot\frac{x-x^{(k)}}{|x-x^{(k)}|}\,\mathrm{d}\sigma-\int_{C^{(k)}_{\varepsilon}}\frac{x}{2}\cdot\frac{x-x^{(k)}}{|x-x^{(k)}|}\,\mathrm{d}\sigma.$
With (4.20) and since $|x|\leqslant 1$ on $C^{(k)}_{\varepsilon}$, we obtain that
$\left|\int_{C^{(k)}_{\varepsilon}}\nabla f_{k^{\prime}}\cdot\vec{n}\,\mathrm{d}\sigma\right|\leqslant\left|C^{(k)}_{\varepsilon}\right|\left(\frac{1}{\rho}+\frac{1}{2}\right).$
Therefore,
(4.25) $\left|-\frac{1}{\sqrt{\pi}}\sum_{\begin{subarray}{c}k^{\prime}=1\\ k^{\prime}\neq k\end{subarray}}^{N}K^{(k^{\prime})}_{\varepsilon}\int_{C^{(k)}_{\varepsilon}}\nabla f_{k^{\prime}}\cdot\vec{n}\,\mathrm{d}\sigma\right|\leqslant\frac{1}{\sqrt{\pi}}\left(\frac{1}{\rho}+\frac{1}{2}\right)\left(\sum_{\begin{subarray}{c}k^{\prime}=1\\ k^{\prime}\neq k\end{subarray}}^{N}K^{(k^{\prime})}_{\varepsilon}\right)r^{(k)}_{\varepsilon,+}(\pi-\theta_{\varepsilon})=\mathrm{O}\left(\overline{K}^{2}_{\varepsilon}\right).$
Gathering the latter inequality with (4.17) and (4.19) finally gives the desired result. ∎
We are now in position to prove Theorem 1.4.
Proof of Theorem 1.4.
Note first that the estimate (1.28) is given by Lemma 4.1. Fix next $k\in\{1,\dots,N\}$. Using Lemmas 4.3 and 4.4, we have, for $\varepsilon$ small enough,
(4.26) $\int_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}}\nabla{\varphi}^{\varepsilon}\cdot\vec{n}\,\mathrm{d}\sigma=\sqrt{\pi}{K}^{(k)}_{\varepsilon}+\mathrm{O}\left(\overline{K}^{2}_{\varepsilon}\right).$
Lemma 4.2 and (4.26) then imply that, for $\varepsilon$ small enough,
(4.27) $\begin{split}\int_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}}\nabla\widetilde{u}_{0}^{\varepsilon}\cdot\vec{n}\,\mathrm{d}\sigma&=\int_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}}\nabla{\varphi}^{\varepsilon}\cdot\vec{n}\,\mathrm{d}\sigma+\int_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}}\nabla(\widetilde{u}_{0}^{\varepsilon}-{\varphi}^{\varepsilon})\cdot\vec{n}\,\mathrm{d}\sigma\\ &=\sqrt{\pi}{K}^{(k)}_{\varepsilon}+\mathrm{O}\left(\overline{K}^{3/2}_{\varepsilon}\right),\end{split}$
which is (1.29). Moreover, using (4.1), (1.28), and (1.29), one obtains that, for $\varepsilon$ small enough,
$\mathbb{P}_{\widetilde{\nu}_{0}^{\varepsilon}}\left[X_{\widetilde{\tau_{\varepsilon}}}\in\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}\right]=\frac{\displaystyle\int_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}_{k}}}\nabla\widetilde{u}_{0}^{\varepsilon}\cdot\vec{n}\,\mathrm{d}\sigma}{\displaystyle\int_{\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}}\nabla\widetilde{u}_{0}^{\varepsilon}\cdot\vec{n}\,\mathrm{d}\sigma}=\frac{{K}^{(k)}_{\varepsilon}}{\overline{K}_{\varepsilon}}+\mathrm{O}\left(\sqrt{\overline{K}_{\varepsilon}}\right),$
which concludes the proof of Theorem 1.4. ∎
Appendix A The Laplacian on $p$-forms
In this appendix, we provide a spectral analysis of the Laplacian on $p$-forms with mixed tangential-normal boundary conditions. For simplicity, we present the results for the Laplacian on $\Omega$ with tangential (resp. normal) boundary conditions on $\Gamma^{\varepsilon}_{\mathbf{D}}$ (resp. $\Gamma^{\varepsilon}_{\mathbf{N}}$). Let us emphasize that the results presented here also hold with $\Omega=\widetilde{\Omega}_{\varepsilon}$, $\Gamma^{\varepsilon}_{\mathbf{D}}=\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}$ and $\Gamma^{\varepsilon}_{\mathbf{N}}=\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}$ as introduced around (1.18) (see [17] for a general setting for these results, on Riemannian Lispchitz manifolds).
We first need to introduce some definitions and notation from Riemannian geometry (we refer to classical textbooks on Riemannian geometry such as [14] for more details, see also e.g. [31, Appendix]). For $p\in\left\{0,1,2\right\}$, one denotes by $\Lambda^{p}C^{\infty}(\overline{\Omega})$ (respectively $\Lambda^{p}C_{\mathrm{c}}^{\infty}(\Omega)$) the space of $p$-forms which are $C^{\infty}$ on $\Omega$ (respectively $C^{\infty}$ with compact support in $\Omega$). Let us moreover introduce the space
(A.1) $\Lambda^{p}C_{0}^{\infty}(\Omega)=\left\{w\in\Lambda^{p}C^{\infty}(\overline{\Omega})\ \middle|\ \textbf{t}w_{|_{\Gamma^{\varepsilon}_{\mathbf{D}}}}=0\hskip 5.69046pt\text{and}\hskip 5.69046pt\textbf{n}w_{|_{\Gamma^{\varepsilon}_{\mathbf{N}}}}=0\right\},$
where t denotes the tangential trace and n the normal trace on forms. One denotes by $\Lambda^{p}H^{m}(\Omega)$ the Sobolev spaces of $p$-forms with regularity index $m$ on $\Omega$, the space $\Lambda^{p}H^{0}(\Omega)$ being also denoted by $\Lambda^{p}L^{2}(\Omega)$. We denote by
$\Lambda H^{m}(\Omega)=\bigoplus_{p=0}^{d}\Lambda^{p}H^{m}(\Omega),\qquad\Lambda C^{\infty}(\Omega)=\bigoplus_{p=0}^{d}\Lambda^{p}C^{\infty}(\Omega),$
and by $d$ the differential on $\Lambda C^{\infty}(\Omega)$:
(A.2) $d^{(p)}:\Lambda^{p}C^{\infty}(\Omega)\to\Lambda^{p+1}C^{\infty}(\Omega).$
Moreover, $d^{\star}$ denotes its formal adjoint with respect to the $L^{2}$-scalar product inherited from the Riemannian structure:
(A.3) $d^{(p),\star}:\Lambda^{p+1}C^{\infty}(\Omega)\to\Lambda^{p}C^{\infty}(\Omega).$
The following proposition is taken from [17, Proposition 4.4].
Proposition A.1.
The unbounded operators $d^{(p)}$ and $d^{(p),\star}$ defined on $\Lambda^{p}L^{2}(\Omega)$ with domains
(A.4) $\mathcal{D}\left(d^{(p)}\right)=\left\{w\in\Lambda^{p}L^{2}({\Omega})\ \middle|\ d^{(p)}w\in\Lambda^{p+1}L^{2}({\Omega}),\,\mathbf{t}w_{|_{\Gamma^{\varepsilon}_{\mathbf{D}}}}=0\right\}$
and
(A.5) $\mathcal{D}\left(d^{(p),\star}\right)=\left\{w\in\Lambda^{p}L^{2}({\Omega})\ \middle|\ d^{(p),\star}w\in\Lambda^{p-1}L^{2}({\Omega}),\,\mathbf{{n}}w_{|_{\Gamma^{\varepsilon}_{\mathbf{N}}}}=0\right\},$
are closed, densely defined, and adjoints of each other in $\Lambda^{p}L^{2}(\Omega)$.
On can check that (see [15, Equation (130)])
[…]
We are now in position to define the Laplacians on $p$-forms with mixed tangential-normal boundary conditions on $\partial\Omega$ (see also [15, p. 89]). The following result is a consequence of [17, Theorem 4.5].
Proposition A.2.
Define on $\Lambda^{p}L^{2}(\Omega)$ the Laplacian on $p$-forms
(A.9) $\Delta^{(p)}=d^{(p-1)}\circ d^{(p),\star}+d^{(p+1),\star}\circ d^{(p)},$
with domain
(A.10) $\begin{split}\mathcal{D}\left(\Delta^{(p)}\right)=\biggl{\{}&w\in\Lambda^{p}L^{2}({\Omega})\ \Big{|}\,\,dw,\,d^{\star}w,\,dd^{\star}w,\,d^{\star}dw\in\Lambda L^{2}({\Omega}),\\ &\qquad\mathbf{t}w_{|_{\Gamma^{\varepsilon}_{\mathbf{D}}}}=0,\,\mathbf{n}w_{|_{\Gamma^{\varepsilon}_{\mathbf{N}}}}=0,\,\mathbf{t}d^{\star}w_{|_{\Gamma^{\varepsilon}_{\mathbf{D}}}}=0,\,\mathbf{n}dw_{|_{\Gamma^{\varepsilon}_{\mathbf{N}}}}=0\biggr{\}}.\end{split}$
This operator is nonnegative selfadjoint in $\Lambda^{p}L^{2}(\Omega)$. In addition, the domain $\mathcal{D}(Q^{(p)})$ of the closed quadratic form $Q^{(p)}$ associated with $\Delta^{(p)}$ is given by
(A.11) $\begin{split}\mathcal{D}\left(Q^{(p)}\right)&=\mathcal{D}\left(d^{(p)}\right)\cap\mathcal{D}\left(d^{(p),\star}\right)\\ &=\biggl{\{}w\in\Lambda^{p}L^{2}({\Omega})\Big{|}\,\,dw,\,d^{\star}w\in\Lambda L^{2}({\Omega}),\,\mathbf{t}w_{|_{\Gamma^{\varepsilon}_{\mathbf{D}}}}=0,\,\mathbf{n}w_{|_{\Gamma^{\varepsilon}_{\mathbf{N}}}}=0\biggr{\}},\end{split}$
and, for $w\in\mathcal{D}(Q^{(p)})$,
(A.12) $Q^{(p)}(w)=\|dw\|^{2}_{L^{2}(\Omega)}+\|d^{\star}w\|^{2}_{L^{2}(\Omega)}.$
In order to make a link with the notation used in the main body of this article, $\mathcal{L}^{\varepsilon}$ is nothing else than the operator $\Delta^{(0)}$, and ~$\widetilde{\mathcal{L}}^{\varepsilon}$ is the operator $\Delta^{(0)}$ with $\Omega=\widetilde{\Omega}_{\varepsilon}$, $\Gamma^{\varepsilon}_{\mathbf{D}}=\widetilde{\Gamma}^{\varepsilon}_{\mathbf{D}}$ and $\Gamma^{\varepsilon}_{\mathbf{N}}=\widetilde{\Gamma}^{\varepsilon}_{\mathbf{N}}$. Notice in particular that in this section, for the ease of notation, we do not explicitly indicate the dependence of the operators $\Delta^{(p)}$ on $\varepsilon$.
The following proposition concerns the commutation property between the spectral projectors of $\Delta^{(p)}$ (for $p\in\left\{0,1,2\right\}$) and the differential and codifferential operators. These are standard results that we recall since they are used in the main body of the article.
Proposition A.3.
Consider the Laplacian on $p$-forms defined in (A.9) and (A.10), as well as the form domain $\mathcal{D}(Q^{(p)})$ defined in (A.11). The differential $d$ and codifferential $d^{\star}$ satisfy the following commutation property: for all $z\in\mathbb{C}\backslash\sigma(\Delta^{(p)})$ and $w\in\mathcal{D}(Q^{(p)})$,
[…]
Consequently, for any $t\in\mathbb{R}_{+}$,
(A.14) $d\circ\pi_{[0,\,t]}\left(\Delta^{(p)}\right)=\pi_{[0,\,t]}\left(\Delta^{(p+1)}\right)\circ d,\qquad d^{\star}\circ\pi_{[0,\,t]}\left(\Delta^{(p)}\right)=\pi_{[0,\,t]}\left(\Delta^{(p-1)}\right)\circ d^{\star}.$
Proof.
Since $\Lambda^{p}C^{\infty}(\Omega)\cap\mathcal{D}(Q^{(p)})$ is dense in $\mathcal{D}(Q^{(p)})$, it is sufficient to consider the case of smooth forms $w\in\mathcal{D}(Q^{(p)})$. For $z\in\mathbb{C}\backslash\sigma(\Delta^{(p)})$, let us introduce
(A.15) $u=\left(z-\Delta^{(p)}\right)^{-1}w.$
Due to the ellipticity of the associated boundary problem, (A.15) implies that $u$ belongs to $\Lambda^{p}C^{\infty}(\Omega)$. Using the fact that $d$ and $d^{\star}$ commute with $\Delta^{(p)}$, we obtain the following relations as differential operators on $\Omega$:
(A.16) $d\left(z-\Delta^{(p)}\right)u=\left(z-\Delta^{(p+1)}\right)du=dw,$
and
(A.17) $d^{\star}\left(z-\Delta^{(p)}\right)u=\left(z-\Delta^{(p+1)}\right)d^{\star}u=d^{\star}w.$
Since $u$ belongs to $\mathcal{D}(\Delta^{(p)})$, it satisfies
(A.18) $\textbf{t}u_{|_{\Gamma^{\varepsilon}_{\mathbf{D}}}}=0,\quad\textbf{n}u_{|_{\Gamma^{\varepsilon}_{\mathbf{N}}}}=0,\quad\textbf{t}d^{\star}u_{|_{\Gamma^{\varepsilon}_{\mathbf{D}}}}=0,\quad\textbf{n}du_{|_{\Gamma^{\varepsilon}_{\mathbf{N}}}}=0.$
Let us check that $du\in\mathcal{D}(\Delta^{(p+1)})$. Using [31, Appendix], we have that t commutes with the differential. Using (A.18), we obtain that $\textbf{t}du=d\textbf{t}u=0$ on $\Gamma^{\varepsilon}_{\mathbf{D}}$. We have that $\textbf{t}d^{\star}(du)=\textbf{t}\Delta^{(p)}u-\textbf{t}dd^{\star}u=z\textbf{t}u-\textbf{t}w-d\textbf{t}d^{\star}u=0$ on $\Gamma^{\varepsilon}_{\mathbf{D}}$, where we used (A.9) and (A.18). We get directly $\textbf{n}du=0$ on $\Gamma^{\varepsilon}_{\mathbf{N}}$ from (A.18) and $\textbf{n}d(du)=0$ from the fact that $d\circ d=0$. Hence $du\in\mathcal{D}(\Delta^{(p+1)})$ and the identity (A.16) yield
$d\left(z-\Delta^{(p)}\right)^{-1}w=du=\left(z-\Delta^{(p+1)}\right)^{-1}dw,$
which proves the first commutation relation. For the second one, it is sufficient to show that $d^{\star}u\in\mathcal{D}(\Delta^{(p-1)})$. Using [31, Appendix], we have that n commutes with the codifferential. With (A.18), we obtain that $\textbf{n}d^{\star}u=d^{\star}\textbf{n}u=0$ on $\Gamma^{\varepsilon}_{\mathbf{N}}$. We have that $\textbf{n}d(d^{\star}u)=\textbf{n}\Delta^{(p)}u-\textbf{n}d^{\star}du=z\textbf{n}u-\textbf{n}w-d^{\star}\textbf{n}du=0$ on $\Gamma^{\varepsilon}_{\mathbf{N}}$, where we used (A.9) and (A.18). We directly get $\textbf{t}d^{\star}u=0$ on $\Gamma^{\varepsilon}_{\mathbf{D}}$ from (A.18) and $\textbf{t}d^{\star}(d^{\star}u)=0$ from the fact that $d^{\star}\circ d^{\star}=0$. Finally, (A.14) follows directly from (A.13) and Stone’s formula. ∎
Acknowledgments
We thank Doug Arnold (University of Minnesota) and Martin Licht (EPFL) for discussions on the appropriate discretization of the eigenvalue problems, and Jean-François Bony (Université de Bordeaux), Dorian Le Peutrec (Nantes Université), Laurent Michel (Université de Bordeaux) and Boris Nectoux (Université Clermont Auvergne) for discussions on the mathematical analysis. This work was funded by the Agence Nationale de la Recherche, under grant ANR-19-CE40-0010-01 (QuAMProcs), and by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (project EMC2, grant agreement No 810367).